### v3

# A Malliavin calculus approach to score functions in diffusion generative models

Abstract
Score-based diffusion generative models have recently emerged as a powerful tool for modelling complex data distributions. These models aim at learning the score function, which defines a map from a known probability distribution to the target data distribution via deterministic or stochastic differential equations (SDEs). The score function is typically estimated from data using a variety of approximation techniques, such as denoising or sliced score matching, Hyvärien’s method, or Schrödinger bridges. In this paper, we derive an exact, closed-form, expression for the score function for a broad class of nonlinear diffusion generative models. Our approach combines modern stochastic analysis tools such as Malliavin derivatives and their adjoint operators (Skorokhod integrals or Malliavin Divergence) with a new Bismut-type formula. The resulting expression for the score function can be written entirely in terms of the first and second variation processes, with all Malliavin derivatives systematically eliminated, thereby enhancing its practical applicability. The theoretical framework presented in this work offers a principled foundation for advancing score estimation methods in generative modelling, enabling the design of new sampling algorithms for complex probability distributions. Our results can be extended to broader classes of stochastic differential equations, opening new directions for the development of score-based diffusion generative models.
60H07,
60H10; 60J60; 68T07,
Score function,
Bismut formula,
Malliavin calculus,
stochastic differential equations,
generative modelling.,
keywords:
[class=MSC]
keywords:
[…]
and
1 Introduction
Score-based diffusion generative models [24, 23] have recently emerged as a powerful tool for modelling complex data distributions in a variety of applications ranging from finance [22] to image synthesis [4, 21]. These models aim to estimate the so-called score function $\nabla_{y}\log p_{t}(y)$, which governs the transformation of a simple reference distribution into the target distribution via a time-reversed nonlinear diffusion process.
Historically, score-based generative models have been built upon diffusion processes described by linear Itô’s stochastic differential equations, i.e., SDEs with drift and diffusion coefficients that are linear in the state variables. Such a linearity assumption ensures analytical tractability and comes with other desirable properties. For example, the Fokker–Planck equation [7, 20] associated with the diffusion process admits a Gaussian stationary distribution and Gaussian transition densities, which have been used in the development of techniques such as Denoising Diffusion Probabilistic Models (DDPM) [9]. While the linearity assumption in diffusion processes has been instrumental in developing the first generation of score-based generative models, it also represents an overly restrictive simplification. Many real-world systems exhibit nonlinear dynamics, where both the drift and diffusion coefficients depend nonlinearly on the state variables. Extending diffusion models to such settings poses substantial challenges. Specifically, unlike linear SDEs, nonlinear SDEs typically lack closed-form expressions for their transition probability densities, making the computation of the score function $\nabla_{y}\log p_{t}(y)$ very challenging.
The purpose of this paper is to bridge this gap by establishing a new rigorous link between score-based diffusion models and Malliavin calculus, a stochastic calculus of variations introduced by Paul Malliavin in the 1970s [11, 12]. Originally devised to investigate hypoelliptic operators and stochastic partial differential equations, Malliavin calculus provides a robust framework for analysing the smoothness of SDE solutions and their densities. By leveraging Malliavin derivatives and a novel Bismut-type formula we derive in this paper, which is distinct from the classical Bismut–Elworthy–Li formula [3, 5, 6] tailored to heat kernels, we derive an explicit (closed-form) representation of the score function for general nonlinear diffusion processes. The resulting formula is expressed entirely in terms of first and second variation processes [10, 2], with all Malliavin derivatives systematically eliminated. In particular, our framework generalises the state‐independent results of [14] to fully nonlinear SDEs with state‐dependent diffusion coefficients. For the numerical experiments of this framework, see the same work [14]. Very recently, other studies have applied Malliavin calculus to diffusion generative models for different ends: one focuses on robust numerical treatment of singular diffusion bridges and conditional control, and additionally derives an explicit yet distinct conditional-score formula via a generalised Tweedie identity [19], and another proposes an extension of diffusion models to infinite‐dimensional diffusion settings using Malliavin calculus [8].
The implications of this new theoretical framework are profound. First, Malliavin calculus provides a powerful tool for overcoming the analytical challenges posed by nonlinear diffusion processes. Second, its rigorous foundations enable the development and analysis of new classes of generative models beyond linear diffusion. From a practical standpoint, this can potentially yield more accurate generative models for applications such as high-fidelity image synthesis, fluid dynamics [25], and protein folding simulations in molecular dynamics [1].
This paper is organised as follows. In Section 2 we state our main theorem expressing the score function $\nabla_{y}\log p_{t}(y)$ in terms of first and second variation processes via a Bismut-type formula. Unlike the Bismut-Elworthy-Li formula, our formula directly targets the score function, leveraging the first and second variation processes to yield a practical closed-form expression. In Section 3 we briefly review basic notions of Malliavin calculus and prove various theorems on covering vector fields. In Section 4 we discuss first and second variation processes in detail. In Section 5 we prove our main theorem, i.e., Theorem 2.1. The proof includes the full derivation of the score function for nonlinear SDEs, along with several useful lemmas that assist in the process. In Section 6 we provide a simplified expression of the score function for nonlinear SDEs with state-independent diffusion coefficients.
2 Problem statement and main result
Consider the $m$-dimensional stochastic differential equation
$dX_{t}=b(t,X_{t})\,dt+\sigma(t,X_{t})\,dB_{t},\qquad X_{0}=x,\qquad 0\leq t\leq T,$ (2.1)
where $X_{t}\in\mathbb{R}^{m}$, $B_{t}\in\mathbb{R}^{d}$ is a standard Brownian motion, $b:[0,T]\times\mathbb{R}^{m}\to\mathbb{R}^{m}$ is the drift coefficient, and $\sigma:[0,T]\times\mathbb{R}^{m}\to\mathbb{R}^{m\times d}$ is the diffusion coefficient, with $\sigma^{l}(t,X_{t})$ denoting its $l$-th column for $l=1,\ldots,d$. The initial condition $x\in\mathbb{R}^{m}$ is deterministic, and $T>0$ is a fixed terminal time. We assume that $b$ and $\sigma$ are $\mathcal{C}^{2}$ functions with bounded derivatives, which ensures the existence of a unique strong solution to (2.1).
In the framework of score-based diffusion models, the SDE (2.1) represents the so-called forward diffusion process, which gradually transforms the initial data distribution into some other distribution over the time interval $[0,T]$. To enable generative sampling, we define the corresponding reverse-time diffusion process, which runs backwards in time from $T$ to 0. Based on the general structure of reverse-time SDEs, this process is given by
$dX_{t}=\left\{b(t,X_{t})-\nabla\cdot\left[\sigma(t,X_{t})\sigma(t,X_{t})^{\top}\right]+\sigma(t,X_{t})\nabla_{X_{t}}\log p_{t}(X_{t})\right\}\,dt+\sigma(t,X_{t})\,d\bar{B}_{t},$ (2.2)
for all $0\leq t\leq T$, where $\nabla\cdot[\sigma(t,X_{t})\sigma(t,X_{t})^{\top}]$ denotes the divergence of the matrix $\sigma(t,X_{t})\sigma(t,X_{t})^{\top}$ (a correction term arising from the Itô interpretation of the SDE), $\nabla_{X_{t}}\log p_{t}(X_{t})$ is the score function, and $d\bar{B}_{t}$ represents the increment of a reverse-time Brownian motion. The reverse-time SDE effectively reverses the forward diffusion by leveraging the score function to guide the process from a noise distribution back to the original data distribution. Our goal is to compute the score function
$\nabla_{y}\log p_{t}(y)=\frac{\nabla_{y}p_{t}(y)}{p_{t}(y)},\quad\partial_{y_{k}}\log p_{t}(y)=\frac{\partial_{y_{k}}p_{t}(y)}{p_{t}(y)},\quad k=1,\ldots,m,$
which is essential for implementing the reverse diffusion process in generative tasks. In particular, we are interested in $\nabla_{y}\log p_{T}(y)$ [^1] For generative tasks, one requires the time-dependent score $\nabla_{y}\log p_{t}(y)$ for each intermediate time $t$, which is obtained by evaluating the integrals discussed in Theorem 2.1 at every such $t$.. To achieve this goal, we leverage Malliavin calculus and a new Bismut-type formula we develop in this paper. As we shall see hereafter, this allows us to express the score in terms of the first variation process $Y_{t}=\partial X_{t}/\partial x$ and the second variation process $Z_{t}=\partial^{2}X_{t}/\partial x^{2}$, which capture the first- and the second-order sensitivity of the solution $X_{t}$ relative to changes in the initial condition $x$. Our main result is the following theorem.
Theorem 2.1 (Skorokhod integral representation theorem for the score function).
Let $X_{t}$ be the solution to the stochastic differential equation
$dX_{t}=b(t,X_{t})\,dt+\sigma(t,X_{t})\,dB_{t},\quad X_{0}=x_{0},$
where $b:[0,T]\times\mathbb{R}^{m}\to\mathbb{R}^{m}$ and $\sigma:[0,T]\times\mathbb{R}^{m}\to\mathbb{R}^{m\times d}$ are sufficiently smooth functions, and $B_{t}$ is a $d$-dimensional Brownian motion. Define the first variation process
$dY_{t}=\partial_{x}b(t,X_{t})Y_{t}\,dt+\sum_{l=1}^{d}\partial_{x}\sigma^{l}(t,X_{t})Y_{t}\,dB_{t}^{l},\quad Y_{0}=I_{m},$
and the Malliavin covariance matrix of $X_{T}$ as
$\gamma_{X_{T}}=\int_{0}^{T}Y_{T}Y_{s}^{-1}\sigma(s,X_{s})\sigma(s,X_{s})^{\top}(Y_{s}^{-1})^{\top}Y_{T}^{\top}\,ds$
which we assume to be invertible almost surely. Furthermore, define the random field $u_{t}(x)$ and the random vector $F_{k}$ as follows [^2] Note that $u_{t}(x)\in\mathbb{R}^{d}$ since $x^{\top}$ is a $1\times m$ row vector, $Y_{t}^{-1}$ is an $m\times m$ matrix, and $\sigma(t,X_{t})$ is an $m\times d$ matrix, resulting in a $1\times d$ row vector. Moreover, $u_{t}(x)$ is adapted to the natural Brownian filtration $\mathcal{F}_{t}$ for each fixed $x$, as it depends only on $Y_{t}^{-1}$ and $\sigma(t,X_{t})$, both of which are adapted processes.
$u_{t}(x)$ $=x^{\top}Y_{t}^{-1}\sigma(t,X_{t}),\quad x\in\mathbb{R}^{m},$
$F_{k}$ $=Y_{T}^{\top}\gamma_{X_{T}}^{-1}e_{k},\quad\text{with}\quad F_{k}^{j}=e_{j}^{\top}Y_{T}^{\top}\gamma_{X_{T}}^{-1}e_{k},$
where $Y_{t}$ is the first variation process, $Y_{T}$ is its value at time $T$, $Y_{t}^{-1}$ its inverse at time $t$, and $e_{k}$ is the $k$-th standard basis vector in $\mathbb{R}^{m}$. Then, following [16], for each $k=1,\dots,m$, the gradient of the log-density satisfies
$\partial_{y_{k}}\log{p_{T}}(y)=-\mathbb{E}\left[\delta(u_{k})\mid X_{T}=y\right],$ (2.3)
where $\delta(u_{k})$ denotes the Skorokhod integral of the process $u_{k}(t)$, i.e., the Malliavin divergence, whose explicit representation is given by
$\delta(u_{k})$ $=\left.\int_{0}^{T}u_{t}(x)\cdot dB_{t}\right|_{x=F_{k}}-\int_{0}^{T}\sum_{j=1}^{m}\bigl{[}Y_{t}^{-1}\sigma(t,X_{t})\bigr{]}_{j}\cdot A_{jk}(t)\,dt$
$\quad+\int_{0}^{T}\sum_{j=1}^{m}\bigl{[}Y_{t}^{-1}\sigma(t,X_{t})\bigr{]}_{j}\cdot B_{jk}(t)\,dt+\int_{0}^{T}\sum_{j=1}^{m}\bigl{[}Y_{t}^{-1}\sigma(t,X_{t})\bigr{]}_{j}\cdot C_{jk}(t)\,dt,$ (2.4)
where
$A_{jk}(t)$ $=e_{j}^{\top}\bigg{[}\sigma(t,X_{t})^{\top}(Y_{t}^{-1})^{\top}Z_{T}^{\top}-\sigma(t,X_{t})^{\top}(Y_{t}^{-1})^{\top}Z_{t}^{\top}(Y_{t}^{-1})^{\top}Y_{T}^{\top}$
$\quad\quad\quad+\left(Y_{T}Y_{t}^{-1}\partial_{x}\sigma(t,X_{t})Y_{t}\right)^{\top}\bigg{]}\gamma_{X_{T}}^{-1}e_{k},$
$B_{jk}(t)$ $=e_{j}^{\top}Y_{T}^{\top}\gamma_{X_{T}}^{-1}\cdot\bigg{[}\int_{0}^{t}I_{1}(t,s)\,ds+\int_{0}^{t}I_{2}(t,s)\,ds\bigg{]}\gamma_{X_{T}}^{-1}e_{k},$
$C_{jk}(t)$ $=e_{j}^{\top}Y_{T}^{\top}\gamma_{X_{T}}^{-1}\cdot\bigg{[}\int_{t}^{T}I_{3}(t,s)\,ds+\int_{t}^{T}I_{4}(t,s)\,ds\bigg{]}\gamma_{X_{T}}^{-1}e_{k},$
with integrands ($p$ and $q$ below denote components)
$I^{p,q}_{1}(t,s)$ $=\bigg{[}\Omega(t)Y_{s}^{-1}\sigma(s,X_{s})\bigg{]}^{p}\cdot[Y_{T}Y_{s}^{-1}\sigma(s,X_{s})]^{q},$
$I^{p,q}_{2}(t,s)$ $=[Y_{T}Y_{s}^{-1}\sigma(s,X_{s})]^{p}\cdot\bigg{[}\Omega(t)Y_{s}^{-1}\sigma(s,X_{s})\bigg{]}^{q},$
$I^{p,q}_{3}(t,s)$ $=\bigg{[}\Omega(t)Y_{s}^{-1}\sigma(s,X_{s})+Y_{T}\Theta(t,s)\bigg{]}^{p}\cdot[Y_{T}Y_{s}^{-1}\sigma(s,X_{s})]^{q},$
$I^{p,q}_{4}(t,s)$ $=[Y_{T}Y_{s}^{-1}\sigma(s,X_{s})]^{p}\cdot\bigg{[}\Omega(t)Y_{s}^{-1}\sigma(s,X_{s})+Y_{T}\Theta(t,s)\bigg{]}^{q},$
and
$\Omega(t)$ $=Z_{T}Y_{t}^{-1}\sigma(t,X_{t})-Y_{T}Y_{t}^{-1}Z_{t}Y_{t}^{-1}\sigma(t,X_{t})+Y_{T}Y_{t}^{-1}\partial_{x}\sigma(t,X_{t})Y_{t},$
$\Theta(t,s)$ $=-Y_{s}^{-1}\bigg{[}Z_{s}Y_{t}^{-1}\sigma(t,X_{t})-Y_{s}Y_{t}^{-1}Z_{t}Y_{t}^{-1}\sigma(t,X_{t})$
$\quad+Y_{s}Y_{t}^{-1}\partial_{x}\sigma(t,X_{t})Y_{t}\bigg{]}Y_{s}^{-1}\sigma(s,X_{s})+Y_{s}^{-1}\partial_{x}\sigma(s,X_{s})\bigg{(}Y_{s}Y_{t}^{-1}\sigma(t,X_{t})\bigg{)}.$
The second variation process appearing in $\Omega(t)$ and $\Theta(t,s)$ is defined as
$dZ_{t}$ $=\left[\partial_{xx}b(t,X_{t})(Y_{t}\otimes Y_{t})+\partial_{x}b(t,X_{t})Z_{t}\right]dt$
$\quad+\sum_{l=1}^{d}\left[\partial_{xx}\sigma^{l}(t,X_{t})(Y_{t}\otimes Y_{t})+\partial_{x}\sigma^{l}(t,X_{t})Z_{t}\right]dB_{t}^{l},$
with initial condition $Z_{0}=0$.
The formula (2.4) for the Skorokhod integral $\delta(u_{k})$ consists of several components that capture the interaction between stochastic processes and their Malliavin derivatives. The first term $\left.\int_{0}^{T}u_{t}(x)\cdot dB_{t}\right|_{x=F_{k}}$ represents the Itô integral from Bismut’s formula, evaluated at the specific point $x=F_{k}$. This term captures the direct influence of the Brownian motion on the process. The remaining components in the expansion of $\sum_{j=1}^{m}\partial_{j}u_{t}(F_{k})\,\cdot\,D_{t}F_{k}^{j}$ are subtracted from the first component and are indexed by the triplet $(j,p,q)\in\{1,\ldots,m\}^{3}$. This indexing arises from extracting matrix components of the inverse Malliavin matrix $\gamma_{X_{T}}^{-1}$ and selecting vector components of $Y_{t}^{-1}\,\sigma(t,X_{t})$. The formula incorporates integrals over different time intervals, i.e., integrals over $s\in[0,t]$ and integrals over $s\in[t,T]$. These appear in the $D_{t}\gamma_{X_{T}}^{p,q}$ expansions and represent how the Malliavin derivative propagates through the system depending on whether the perturbation time $t$ occurs before or after the reference time $s$.
Note that Theorem 2.1 provides an explicit representation of $\nabla_{y}\log p_{t}(y)$ that avoids abstract Malliavin derivatives, and therefore it is computationally practical for applications in score-based generative modelling.
3 Malliavin calculus and covering vector fields
Malliavin calculus enables us to analyse the regularity of the random process $X_{T}$ and compute derivatives of its density. It also allows us to represent score functions via the Bismut-type formula [^3] The negative sign arises from the integration-by-parts formula in Malliavin calculus, and the expectation is conditioned on $X_{T}=y$, reflecting the evaluation of the density gradient at a specific point.
$\partial_{y_{k}}\log p_{T}(y)=-\mathbb{E}\left[\delta(u_{k})\mid X_{T}=y\right],\qquad k=1,\ldots,m$ (3.1)
where $\delta(u_{k})$ is the Skorokhod integral and $u_{k}=\{u_{k}(t):0\leq t\leq T\}$, with $u_{k}(t)\in\mathbb{R}^{d}$, is a process called the covering vector field, [13, 16]. The covering vector field $u_{k}(t)$ must satisfy
$\langle DX_{T}^{i},u_{k}\rangle_{H}=\delta_{ik},\quad i,k=1,\ldots,m,$ (3.2)
where $\delta_{ik}$ is the Kronecker delta, $DX_{T}^{i}=\{D_{t}X_{T}^{i}:0\leq t\leq T\}$ is the Malliavin derivative of $X_{T}^{i}$, a process in the Hilbert space $H=L^{2}\left([0,T],\mathbb{R}^{d}\right)$,
$\langle f,g\rangle_{H}=\int_{0}^{T}f(t)\cdot g(t)\,dt$
is the inner product in $H$, and $\cdot$ denotes the standard dot product in $\mathbb{R}^{d}$. Condition (3.2) ensures that $u_{k}(t)$ “covers” the $k$-th direction in the Malliavin sense, which will allow us to isolate $\partial_{y_{k}}p(y)$. Define the Malliavin covariance matrix $\gamma_{X_{T}}\in\mathbb{R}^{m\times m}$ with entries
$\gamma_{X_{T}}^{i,j}=\langle DX_{T}^{i},DX_{T}^{j}\rangle_{H},$ (3.3)
where $D_{t}X_{T}^{i}\in\mathbb{R}^{d}$ is the Malliavin derivative of $X_{T}^{i}$ at time $t$. This matrix measures the “randomness” induced by the Brownian motion up to time $T$. We assume $\gamma_{X_{T}}$ is invertible almost surely, so $\gamma_{X_{T}}^{-1}$ exists, which is necessary for $p(y)$ to be smooth. Hereafter we show that for each $k=1,\ldots,m$, the process
$u_{k}(t)=\sum_{j=1}^{m}\left(\gamma_{X_{T}}^{-1}\right)_{k,j}D_{t}X_{T}^{j},$ (3.4)
where $\left(\gamma_{X_{T}}^{-1}\right)_{k,j}$ is the $(k,j)$-th element of the inverse Malliavin matrix $\gamma_{X_{T}}^{-1}$, satisfies the covering condition
$\langle DX_{T}^{i},u_{k}\rangle_{H}=\delta_{ik}\quad\text{for all}\quad i=1,\ldots,m.$
In the remainder of this section, we state and prove several important properties of the chosen covering vector field (3.4). Although some of these results may already be known or established in prior works, we include proofs here for the reader’s convenience and to ensure the exposition remains self-contained, for more details, see [10, 15].
We begin with a theorem establishing the existence and uniqueness of a covering vector field $u_{k}(t)$ for an $\mathbb{R}^{m}$-valued random variable $X_{T}=(X_{T}^{1},\ldots,X_{T}^{m})$, defined on a probability space $(\Omega,\mathcal{F},\mathbb{P})$ with a $d$-dimensional Brownian motion $B=\{B_{t}\}_{t\in[0,T]}$. The theorem assumes that $X_{T}$ lies in $\mathbb{D}^{2,2}$, the space of twice Malliavin differentiable random variables relative to the Gaussian structure of $B$, and that its Malliavin covariance matrix $\gamma_{X_{T}}$, with entries $\gamma_{X_{T}}^{i,j}=\langle DX_{T}^{i},DX_{T}^{j}\rangle_{H}$ in the Cameron–Martin space $H=L^{2}([0,T];\mathbb{R}^{d})$, is invertible almost surely and satisfies $\mathbb{E}[\lambda_{\min}(\gamma_{X_{T}})^{-p}]<\infty$ for some $p>2(m+1)$. Under these conditions, for each $k=1,\ldots,m$, it can be shown that there exists a unique energy-minimising process $u_{k}(t)\in L^{2}([0,T]\times\Omega;\mathbb{R}^{d})$ in the domain of the Skorokhod integral $\mathrm{Dom}(\delta)$, satisfying $\langle DX_{T}^{i},u_{k}\rangle_{H}=\delta_{ik}$, where $\delta_{ik}$ is the Kronecker delta.
Theorem 3.1 (Existence and uniqueness of the covering vector field).
Let $X_{T}=(X_{T}^{1},\ldots,X_{T}^{m})$ be an $\mathbb{R}^{m}$-valued random variable defined on a probability space $(\Omega,\mathcal{F},\mathbb{P})$, equipped with a Brownian motion $B=\{B_{t}\}_{t\in[0,T]}$ taking values in $\mathbb{R}^{d}$. Suppose $X_{T}\in\mathbb{D}^{2,2}$. Let $\gamma_{X_{T}}$ be the Malliavin covariance matrix of $X_{T}$ defined in (3.3). Assume $\gamma_{X_{T}}$ is invertible almost surely with $\mathbb{E}[\lambda_{\min}(\gamma_{X_{T}})^{-p}]<\infty$ for some $p>2(m+1)$.
Then, for each $k=1,\ldots,m$, there exists a unique process $u_{k}=\{u_{k}(t)\}_{t\in[0,T]}$ in $\text{Dom}(\delta)\cap L^{2}([0,T]\times\Omega,\mathbb{R}^{d})$ minimising $\mathbb{E}[\int_{0}^{T}\|u(t)\|^{2}\,dt]$ such that
$\langle DX_{T}^{i},u_{k}\rangle_{H}=\delta_{ik},\quad i=1,\ldots,m.$ (3.5)
Proof.
Set $H=L^{2}([0,T];\mathbb{R}^{d})$, $\gamma=\gamma_{X_{T}}$, and define the covering field by
$u_{k}(t)=\sum_{j=1}^{m}\left(\gamma^{-1}\right)_{jk}D_{t}X_{T}^{j}.$
Define the operator $T:H\to\mathbb{R}^{m}$ by $Th=(\langle DX_{T}^{i},h\rangle_{H})_{i=1}^{m}$. Its adjoint satisfies $T^{*}\mathbf{c}=\sum_{i}c_{i}DX_{T}^{i}$, hence $TT^{*}=\gamma$. Since $TT^{*}=\gamma$, we have
$u_{k}=T^{*}\gamma^{-1}\mathbf{e}_{k},$
which compresses verification and minimality into a standard Moore-Penrose argument.
Let $\mathbf{v}_{k}=(\left(\gamma^{-1}\right)_{1k},\ldots,\left(\gamma^{-1}\right)_{mk})^{\top}$. Then
$\|u_{k}(t)\|^{2}$ $=\sum_{j,\ell}\left(\gamma^{-1}\right)_{jk}\left(\gamma^{-1}\right)_{\ell k}(D_{t}X_{T}^{j}\cdot D_{t}X_{T}^{\ell})\quad\text{(Euclidean inner product in $\mathbb{R}^{d}$)},$ (3.6)
$\int_{0}^{T}\|u_{k}(t)\|^{2}\,dt$ $=\mathbf{v}_{k}^{\top}\gamma\mathbf{v}_{k}.$ (3.7)
Since $\gamma\mathbf{v}_{k}=\mathbf{e}_{k}$, we have $\mathbf{v}_{k}^{\top}\gamma\mathbf{v}_{k}=\left(\gamma^{-1}\right)_{kk}$. Thus
$\mathbb{E}\left[\int_{0}^{T}\|u_{k}(t)\|^{2}\,dt\right]=\mathbb{E}[\left(\gamma^{-1}\right)_{kk}]\leq\mathbb{E}[\lambda_{\min}(\gamma)^{-1}].$
By Hölder’s inequality with conjugate exponents $p$ and p/(p1){p}/{(p-1)},
$\mathbb{E}\!\left[\lambda_{\min}(\gamma)^{-1}\right]\leq\Big{(}\mathbb{E}\!\left[\lambda_{\min}(\gamma)^{-p}\right]\Big{)}^{1/p}<\infty.$
Suppose $v_{k}$ also satisfies (3.5). Then $w=u_{k}-v_{k}\in\ker T$. Since $u_{k}\in\text{range}(T^{*})$ by construction and $\ker T=\text{range}(T^{*})^{\perp}$, we have $\langle u_{k},w\rangle_{H}=0$. Therefore,
$\|v_{k}\|^{2}_{L^{2}}=\|u_{k}-w\|^{2}_{L^{2}}=\|u_{k}\|^{2}_{L^{2}}+\|w\|^{2}_{L^{2}}\geq\|u_{k}\|^{2}_{L^{2}},$
with equality only if $w=0$. The minimality condition forces $w=0$, establishing uniqueness. Since $X_{T}\in\mathbb{D}^{2,2}$, each $\gamma^{ij}\in\mathbb{D}^{1,2}$. The cofactors of $\gamma$ inherit this regularity, and Cramer’s formula gives
$\left(\gamma^{-1}\right)_{jk}=\frac{\text{adj}(\gamma)_{jk}}{\det\gamma}.$
The moment bound $\mathbb{E}[\lambda_{\min}(\gamma)^{-p}]<\infty$ with $p>2(m+1)$ ensures the required integrability via the Malliavin chain rule after localisation (truncate $\det\gamma$ away from 0 and pass to the limit). Since the inverse matrix entries satisfy $\left(\gamma^{-1}\right)_{jk}\in\mathbb{D}^{1,2}$ and $D_{t}X_{T}^{j}\in\mathbb{D}^{1,2}(H)$, the product rule yields $u_{k}\in\mathbb{D}^{1,2}(H)\subset\text{Dom}(\delta)$.
The next theorem concerns the regularity of the covering vector field $u_{k}(t)$ for a stochastic process $\bigl{(}X_{t}\bigr{)}_{t\in[0,T]}$ adapted to a Brownian filtration. It assumes that the terminal value $X_{T}$ lies in $\mathbb{D}^{2,2}$, that the Malliavin covariance matrix $\gamma_{X_{T}}$ is almost surely invertible, and it defines $u_{k}(t)$ as in (3.4). Additional regularity conditions are: uniform ellipticity of $\gamma_{X_{T}}$ (i.e. $\gamma_{X_{T}}\geq\lambda I$ for some $\lambda>0$); and, when $X_{t}$ solves the SDE $\mathrm{d}X_{t}=b(t,X_{t})\,\mathrm{d}t+\sigma(t,X_{t})\,\mathrm{d}B_{t}$, the coefficients $b$ and $\sigma$ are $\mathcal{C}^{2}$ with all derivatives up to second order bounded. Furthermore, under the uniform ellipticity bound one automatically has $\mathbb{E}\bigl{[}\|\gamma_{X_{T}}^{-1}\|^{\,q}\bigr{]}<\infty$ for every $q>0$. Under these hypotheses the theorem shows that each $u_{k}(t)$ belongs to $\mathrm{Dom}\,(\delta)$ and that its Skorokhod integral satisfies $\delta(u_{k})\in L^{p}(\Omega)$ for some $p>1$, thereby confirming the desired regularity of the vector field.
Theorem 3.2 (Regularity of the covering vector field).
Let $X=(X_{t})_{t\in[0,T]}$ be the unique solution of
$\mathrm{d}X_{t}=b(t,X_{t})\,\mathrm{d}t+\sigma(t,X_{t})\,\mathrm{d}B_{t},$
where $b,\sigma\in\mathcal{C}^{2}$ with all mixed derivatives up to order two bounded. Suppose further that
  1. (i)
    $X_{T}\in\mathbb{D}^{2,2}$, and its Malliavin covariance matrix (3.3) is a.s. invertible with $\gamma_{X_{T}}\geq\lambda I_{m}$ for some $\lambda>0$.
  2. (ii)
    Define, for each $k=1,\dots,m$, the covering vector field defined at (3.4)
Then each $u_{k}$ belongs to $\mathrm{Dom}\,(\delta)$, and its Skorokhod integral $\delta(u_{k})$ is well-defined with $\delta(u_{k})\in L^{p}(\Omega)$ for some $p>1$.
Proof.
The Skorokhod integral $\delta$ is the adjoint of the Malliavin derivative operator. A sufficient condition for $u\in\text{Dom}(\delta)$ is that $u\in\mathbb{D}^{1,2}(H)$ with
$\|u\|_{\mathbb{D}^{1,2}(H)}^{2}=\mathbb{E}\left[\int_{0}^{T}|u(t)|^{2}\,dt\right]+\mathbb{E}\left[\int_{0}^{T}\int_{0}^{T}|D_{s}u(t)|^{2}\,ds\,dt\right]<\infty.$
We verify that $u_{k}\in\mathbb{D}^{1,2}(H)$. The Malliavin derivative of $u_{k}(t)$ is
$D_{s}u_{k}(t)=\sum_{j=1}^{m}\left[D_{s}\left((\gamma_{X_{T}}^{-1})_{jk}\right)D_{t}X_{T}^{j}+(\gamma_{X_{T}}^{-1})_{jk}D_{s}D_{t}X_{T}^{j}\right],$
where
$D_{s}(\gamma_{X_{T}}^{-1})_{jk}=-\sum_{p,q=1}^{m}(\gamma_{X_{T}}^{-1})_{jp}(D_{s}(\gamma_{X_{T}})_{pq})(\gamma_{X_{T}}^{-1})_{qk},$
and
$D_{s}(\gamma_{X_{T}})_{pq}=\int_{0}^{T}\left[D_{s}D_{t}X_{T}^{p}\cdot D_{t}X_{T}^{q}+D_{t}X_{T}^{p}\cdot D_{s}D_{t}X_{T}^{q}\right]dt.$
Since $X_{T}\in\mathbb{D}^{2,2}$ and $\|\gamma_{X_{T}}^{-1}\|\leq\lambda^{-1}$, we have $D_{s}(\gamma_{X_{T}}^{-1})_{jk}\in L^{2}(\Omega\times[0,T])$. We need to verify the following two conditions
  1. a)
    $\mathbb{E}\left[\int_{0}^{T}|u_{k}(t)|^{2}\,dt\right]<\infty$,
  2. b)
    $\mathbb{E}\left[\int_{0}^{T}\int_{0}^{T}|D_{s}u_{k}(t)|^{2}\,ds\,dt\right]<\infty$.
For the first condition
$\int_{0}^{T}|u_{k}(t)|^{2}\,dt\leq\|\gamma_{X_{T}}^{-1}\|^{2}\sum_{j=1}^{m}\int_{0}^{T}|D_{t}X_{T}^{j}|^{2}\,dt=\left\|\gamma_{X_{T}}^{-1}\right\|^{2}\operatorname{Tr}\left(\gamma_{X_{T}}\right).$
Since $\|\gamma_{X_{T}}^{-1}\|\leq\lambda^{-1}$ and $\mathbb{E}\left[\operatorname{Tr}\gamma_{X_{T}}\right]<\infty$ (as $X_{T}\in\mathbb{D}^{1,2}$), we obtain
$\mathbb{E}\left[\int_{0}^{T}|u_{k}(t)|^{2}\,dt\right]\leq\lambda^{-2}\mathbb{E}\left[\operatorname{Tr}\gamma_{X_{T}}\right]<\infty.$
For the second condition
$|D_{s}u_{k}(t)|^{2}\leq 2\left|\sum_{j=1}^{m}D_{s}(\gamma_{X_{T}}^{-1})_{jk}D_{t}X_{T}^{j}\right|^{2}+2\left|\sum_{j=1}^{m}(\gamma_{X_{T}}^{-1})_{jk}D_{s}D_{t}X_{T}^{j}\right|^{2}.$ (3.8)
Using Cauchy-Schwarz and integrating
$\mathbb{E}\left[\int_{0}^{T}\int_{0}^{T}|D_{s}u_{k}(t)|^{2}\,ds\,dt\right]$ $\leq 2m\sum_{j=1}^{m}\mathbb{E}\bigg{[}\int_{0}^{T}\int_{0}^{T}\Big{(}|D_{s}(\gamma_{X_{T}}^{-1})_{jk}|^{2}|D_{t}X_{T}^{j}|^{2}$
$\quad+|(\gamma_{X_{T}}^{-1})_{jk}|^{2}|D_{s}D_{t}X_{T}^{j}|^{2}\Big{)}ds\,dt\bigg{]}.$
Since $|D_{s}(\gamma_{X_{T}}^{-1})_{jk}|^{2}\leq C\|D_{s}\gamma_{X_{T}}\|^{2}\in L^{2}(\Omega\times[0,T])$, $|D_{t}X_{T}^{j}|^{2}\in L^{2}(\Omega\times[0,T])$, and $|(\gamma_{X_{T}}^{-1})_{jk}|^{2}\leq\lambda^{-2}$, while $X_{T}\in\mathbb{D}^{2,2}$ ensures $\mathbb{E}\left[\int_{0}^{T}\int_{0}^{T}|D_{s}D_{t}X_{T}^{j}|^{2}\,ds\,dt\right]<\infty$, the second condition holds. Therefore, $u_{k}\in\mathbb{D}^{1,2}(H)\subseteq\text{Dom}(\delta)$ with $\delta(u_{k})\in L^{2}(\Omega)$. For $p=2$, we verify $u_{k}\in\mathbb{D}^{1,2}(H)$ directly
$\mathbb{E}\left[\int_{0}^{T}|u_{k}(t)|^{2}dt\right]\leq\lambda^{-2}\mathbb{E}\left[\operatorname{Tr}(\gamma_{X_{T}})\right]<\infty,$
and similarly for the Malliavin derivative term. Since the Skorokhod integral $\delta:\mathbb{D}^{1,2}(H)\to L^{2}(\Omega)$ is continuous, we have $\delta(u_{k})\in L^{2}(\Omega)\subseteq L^{p}(\Omega)$ for $p=2>1$.
In the next theorem we establish the $L^{2}$-continuity of the covering vector field associated with a stochastic process $(X_{t})_{t\in[0,T]}$, adapted to a Brownian filtration. The process is assumed to have a terminal value $X_{T}$ in the Malliavin space $\mathbb{D}^{2,2}$, with an invertible Malliavin covariance matrix $\gamma_{X_{T}}$. The covering vector field $u_{k}(t)$ is defined using $\gamma_{X_{T}}^{-1}$ and the Malliavin derivatives of $X_{T}$. Additionally, $X_{t}$ satisfies an SDE with $\mathcal{C}^{2}$ coefficients $b$ and $\sigma$ having bounded derivatives up to second order. Under uniform ellipticity of $\gamma_{X_{T}}$, the theorem proves that $t\mapsto u_{k}(t)$ is continuous into $L^{2}(\Omega)$, highlighting the regularity of the vector field in the stochastic setting.
Theorem 3.3 ($L^{2}$–continuity of the covering vector field).
Let $(X_{t})_{t\in[0,T]}$ be the unique solution of
$\mathrm{d}X_{t}\;=\;b(t,X_{t})\,\mathrm{d}t\;+\;\sigma(t,X_{t})\,\mathrm{d}B_{t},\quad X_{0}=x_{0},$
where $b,\sigma\in\mathcal{C}^{2}\bigl{(}[0,T]\times\mathbb{R}^{m}\bigr{)}$ have all first and second derivatives bounded, and $\sigma$ satisfies the linear growth bound $\|\sigma(t,x)\|\leq C(1+\|x\|)$. Suppose additionally that
  1. (i)
    $X_{T}\in\mathbb{D}^{2,2}$ and its Malliavin covariance (3.3) is a.s. invertible with $\gamma_{X_{T}}\geq\lambda I_{m}$ for some $\lambda>0$;
  2. (ii)
    for each $t\in[0,T]$, the mapping $x\mapsto X_{t}^{x}$ is a random diffeomorphism of class $\mathcal{C}^{2}$ almost surely, and for some $p\geq 8$,
    $\mathbb{E}\left[\sup_{u\in[0,T]}\|Y_{u}\|^{p}+\sup_{u\in[0,T]}\|Y_{u}^{-1}\|^{p}\right]<\infty.$
Define for each $k=1,\dots,m$ the covering field (3.4). Then $t\mapsto u_{k}(t)$ is continuous as a map $0,T]\to L^{2}(\Omega)$; equivalently
$\lim_{s\to t}\mathbb{E}\bigl{[}|u_{k}(s)-u_{k}(t)|^{2}\bigr{]}\;=\;0\quad\forall\,t\in[0,T].$
Proof.
Since $b,\sigma\in\mathcal{C}^{2}$ with bounded derivatives, $X_{t}\in\mathbb{D}^{2,2}$. The Malliavin derivative $D_{t}X_{T}$ is given by
$D_{t}X_{T}=Y_{T}Y_{t}^{-1}\sigma(t,X_{t})\mathbf{1}_{\{t\leq T\}},$
where $Y_{t}=\partial X_{t}/\partial x$ is the first variation process satisfying
$dY_{t}=\partial_{x}b(t,X_{t})Y_{t}\,dt+\sum_{l=1}^{d}\partial_{x}\sigma_{l}(t,X_{t})Y_{t}\,dB_{t}^{l},\quad Y_{0}=I_{m}.$
The process $Y_{t}$ is continuous and invertible a.s. with $\mathbb{E}[\|Y_{T}\|^{2}]<\infty$. Note explicitly that $t\mapsto Y_{t}^{-1}$ is a.s. continuous since the inverse map is continuous on $\mathrm{GL}_{m}$ and $t\mapsto Y_{t}$ is a.s. continuous. Also, under our global Lipschitz/linear-growth hypotheses,
$\mathbb{E}\left[\sup_{u\in[0,T]}\|X_{u}\|^{p}\right]<\infty\quad\text{for some }p\geq 8.$
For $s,t\leq T$, we have
$D_{s}X_{T}-D_{t}X_{T}=Y_{T}\left(Y_{s}^{-1}\sigma(s,X_{s})-Y_{t}^{-1}\sigma(t,X_{t})\right).$
Since $Y_{t}^{-1}\sigma(t,X_{t})$ is continuous in $t$ (by continuity of $X_{t}$, $Y_{t}$, and $\mathcal{C}^{2}$-regularity of $\sigma$), and
$\|Y_{s}^{-1}\sigma(s,X_{s})\|$ $\leq C\left(1+\sup_{u\leq T}\|X_{u}\|\right)\sup_{u\leq T}\|Y_{u}^{-1}\|,$
$|D_{s}X_{T}-D_{t}X_{T}|$ $\leq\|Y_{T}\|\|Y_{s}^{-1}\sigma(s,X_{s})-Y_{t}^{-1}\sigma(t,X_{t})\|$
$\leq 2C\|Y_{T}\|\left(1+\sup_{u\leq T}\|X_{u}\|\right)\sup_{u\leq T}\|Y_{u}^{-1}\|,$
hence
$|D_{s}X_{T}-D_{t}X_{T}|^{2}\leq C^{\prime}\|Y_{T}\|^{2}\left(1+\sup_{u\leq T}\|X_{u}\|\right)^{2}\left(\sup_{u\leq T}\|Y_{u}^{-1}\|\right)^{2}.$
By (ii) with $p\geq 8$ and the standard moment bound $\mathbb{E}[\sup_{u\leq T}\|X_{u}\|^{p}]<\infty$, Hölder with exponents $(2,4,4)$ gives the RHS integrable, so dominated convergence theorem applies to yield $\mathbb{E}[|D_{s}X_{T}-D_{t}X_{T}|^{2}]\to 0$. By assumption, $\gamma_{X_{T}}\geq\lambda I_{m}$ a.s., so $\|\gamma_{X_{T}}^{-1}\|\leq\lambda^{-1}$. For the covering field,
$u_{k}(s)-u_{k}(t)=\sum_{j=1}^{m}(\gamma_{X_{T}}^{-1})_{jk}(D_{s}X_{T}^{j}-D_{t}X_{T}^{j}),$
hence by Cauchy-Schwarz,
$\mathbb{E}\left[|u_{k}(s)-u_{k}(t)|^{2}\right]$ $\leq\mathbb{E}\left[\|\gamma_{X_{T}}^{-1}e_{k}\|^{2}\sum_{j=1}^{m}|D_{s}X_{T}^{j}-D_{t}X_{T}^{j}|^{2}\right]$
$\leq\lambda^{-2}\sum_{j=1}^{m}\mathbb{E}\left[|D_{s}X_{T}^{j}-D_{t}X_{T}^{j}|^{2}\right].$
Since each term converges to zero as $s\to t$, we obtain $\lim_{s\to t}\mathbb{E}[|u_{k}(s)-u_{k}(t)|^{2}]=0$.
Next we present a stability theorem for covering vector fields associated with stochastic processes under controlled perturbations. The theorem examines a sequence of stochastic processes $(X_{t}^{n})_{t\in[0,T]}$ and a limiting process $(X_{t})_{t\in[0,T]}$, described by stochastic differential equations driven by the same Brownian motion, with coefficients $b_{n},\sigma_{n}$ and $b,\sigma$, respectively, where these coefficients satisfy smoothness and boundedness conditions (they are $\mathcal{C}^{2}$ with first and second derivatives bounded uniformly in $n$) and converge locally uniformly together with their spatial derivatives (i.e., $C^{1}_{\mathrm{loc}}$-convergence). Both processes are adapted to this Brownian filtration, their terminal values $X_{T}^{n}$ and $X_{T}$ lie in the Malliavin Sobolev space $\mathbb{D}^{1,2}$, and their Malliavin covariance matrices are almost surely invertible with a uniform ellipticity bound; moreover, $\sup_{n}\mathbb{E}\!\left[\|(\gamma^{n})^{-1}\|^{2}\right]<\infty$ and $\mathbb{E}\!\left[\|\gamma^{-1}\|^{2}\right]<\infty$. We also assume uniform moment bounds for the first-variation processes $Y_{t}^{n}=\partial_{x}X_{t}^{n}$ and $(Y_{t}^{n})^{-1}$, and uniform sup-in-time $L^{2}$ bounds for $X^{n}$ and $X$. In addition, we impose a uniform $L^{4}$ Malliavin control, i.e., $\sup_{n}\mathbb{E}\!\int_{0}^{T}\|D_{t}X_{T}^{n}\|^{4}\,dt<\infty$ and $\mathbb{E}\!\int_{0}^{T}\|D_{t}X_{T}\|^{4}\,dt<\infty$ (equivalently, $X_{T}^{n},X_{T}\in\mathbb{D}^{1,4}$). The theorem proves that the covering vector fields $u_{k}^{n}(t)$, defined using the inverse Malliavin covariance matrices and Malliavin derivatives, converge to $u_{k}(t)$ in the $L^{2}([0,T]\times\Omega)$ norm as $n\to\infty$. This result highlights the robustness of the covering vector field construction against perturbations in the SDE dynamics.
Theorem 3.4 (Stability under perturbations).
Let $X^{n}$ and $X$ be the unique strong solutions on $[0,T]$ of
$\mathrm{d}X^{n}_{t}=b_{n}(t,X^{n}_{t})\,\mathrm{d}t+\sigma_{n}(t,X^{n}_{t})\,\mathrm{d}B_{t},\qquad\mathrm{d}X_{t}=b(t,X_{t})\,\mathrm{d}t+\sigma(t,X_{t})\,\mathrm{d}B_{t},$
driven by the same dd–dimensional Brownian motion $B$. Assume:
  1. (i)
    $b_{n},b,\sigma_{n},\sigma\in C^{2}([0,T]\times\mathbb{R}^{m})$ with all first and second derivatives bounded uniformly in $n$, and $(b_{n},\sigma_{n})\to(b,\sigma)$ locally uniformly, and $\partial_{x}b_{n}\to\partial_{x}b,\partial_{x}\sigma_{n}\to\partial_{x}\sigma$ locally uniformly (i.e. $C^{1}_{\text{loc}}$-convergence).
  2. (ii)
    For each $n$ and for $X$, the terminal value lies in $\mathbb{D}^{1,2}$, and their Malliavin covariance matrices
    $\gamma^{n}\;=\;\int_{0}^{T}D_{t}X^{n}_{T}\,(D_{t}X^{n}_{T})^{\top}\,dt,\quad\gamma\;=\;\int_{0}^{T}D_{t}X_{T}\,(D_{t}X_{T})^{\top}\,dt$
    satisfy $\gamma^{n},\gamma\geq\lambda I_{m}>0$ a.s. for some $\lambda>0$, together with $\sup_{n}\mathbb{E}[\|(\gamma^{n})^{-1}\|^{2}]<\infty$ and $\mathbb{E}[\|\gamma^{-1}\|^{2}]<\infty$.
  3. (iii)
    The first variation processes $Y^{n}_{t}=\partial_{x}X^{n}_{t}$ satisfy $\sup_{n,t}\mathbb{E}[\|Y^{n}_{t}\|^{2}+\|(Y^{n}_{t})^{-1}\|^{2}]<\infty$, and $\sup_{n}\mathbb{E}\big{[}\sup_{t\in[0,T]}\|X^{n}_{t}\|^{2}\big{]}<\infty$ and $\mathbb{E}\big{[}\sup_{t\in[0,T]}\|X_{t}\|^{2}\big{]}<\infty$.
  4. (iv)
    (Uniform $L^{4}$ Malliavin control) $\sup_{n}\mathbb{E}\!\int_{0}^{T}\|D_{t}X_{T}^{n}\|^{4}\,dt<\infty\quad\text{and}\quad\mathbb{E}\!\int_{0}^{T}\|D_{t}X_{T}\|^{4}\,dt<\infty.$
Define, for each $k=1,\dots,m$,
$u^{n}_{k}(t)\;=\;\sum_{j=1}^{m}(\gamma^{n})^{-1}_{jk}\,D_{t}X_{T}^{n,j},\quad u_{k}(t)\;=\;\sum_{j=1}^{m}\gamma^{-1}_{jk}\,D_{t}X_{T}^{j}.$
Then for every $k$,
$\lim_{n\to\infty}\mathbb{E}\!\Bigl{[}\!\int_{0}^{T}\bigl{|}u^{n}_{k}(t)-u_{k}(t)\bigr{|}^{2}\,dt\Bigr{]}\;=\;0.$
Proof.
The proof proceeds in four steps: establishing convergence of $X_{T}^{n}\to X_{T}$, $D_{t}X_{T}^{n}\to D_{t}X_{T}$, $(\gamma^{n})^{-1}\to\gamma^{-1}$, and finally $u_{k}^{n}\to u_{k}$. To show $\mathbb{E}[|X_{T}^{n}-X_{T}|^{2}]\to 0$, we write
$X_{t}^{n}-X_{t}=\int_{0}^{t}[b_{n}(s,X_{s}^{n})-b(s,X_{s})]ds+\int_{0}^{t}[\sigma_{n}(s,X_{s}^{n})-\sigma(s,X_{s})]dB_{s}.$
Using $(a+b)^{2}\leq 2a^{2}+2b^{2}$ and taking suprema:
$\mathbb{E}\left[\sup_{t\in[0,T]}|X_{t}^{n}-X_{t}|^{2}\right]$ $\leq 2\mathbb{E}\left[\sup_{t\in[0,T]}\left|\int_{0}^{t}[b_{n}(s,X_{s}^{n})-b(s,X_{s})]ds\right|^{2}\right]$
$\quad+2\mathbb{E}\left[\sup_{t\in[0,T]}\left|\int_{0}^{t}[\sigma_{n}(s,X_{s}^{n})-\sigma(s,X_{s})]dB_{s}\right|^{2}\right].$
For the drift term, by Cauchy-Schwarz
$\left(\int_{0}^{T}|b_{n}(s,X_{s}^{n})-b(s,X_{s})|ds\right)^{2}\leq T\int_{0}^{T}|b_{n}(s,X_{s}^{n})-b(s,X_{s})|^{2}ds.$
For the diffusion term, by the Burkholder-Davis-Gundy inequality
$\mathbb{E}\left[\sup_{t\in[0,T]}\left|\int_{0}^{t}[\sigma_{n}(s,X_{s}^{n})-\sigma(s,X_{s})]dB_{s}\right|^{2}\right]$ $\leq C_{\mathrm{BDG}}\mathbb{E}\left[\int_{0}^{T}\|\sigma_{n}(s,X_{s}^{n})-\sigma(s,X_{s})\|^{2}ds\right].$
Since the convergence $(b_{n},\sigma_{n})\to(b,\sigma)$ is locally uniform, we employ a localisation argument. Define the stopping time $\tau_{R}=\inf\{t:\max(\|X_{t}\|,\|X_{t}^{n}\|)>R\}$. On $0,\tau_{R}]$, the processes remain in $B_{R}$, where the convergence is uniform. Decomposing the differences using this uniform convergence on $B_{R}$ and Lipschitz continuity
$|b_{n}(s,X_{s}^{n})-b(s,X_{s})|^{2}$ $\leq 2|b_{n}(s,X_{s}^{n})-b(s,X_{s}^{n})|^{2}+2|b(s,X_{s}^{n})-b(s,X_{s})|^{2}$
$\leq 2\sup_{(t,x)\in[0,T]\times B_{R}}|b_{n}(t,x)-b(t,x)|^{2}+2C^{2}|X_{s}^{n}-X_{s}|^{2},$
and similarly for $\sigma$. This yields
$\mathbb{E}\left[\sup_{t\in[0,\tau_{R}]}|X_{t}^{n}-X_{t}|^{2}\right]$ $\leq 4T^{2}\sup_{(t,x)\in[0,T]\times B_{R}}|b_{n}(t,x)-b(t,x)|^{2}$
$\quad+4C_{\mathrm{BDG}}T\sup_{(t,x)\in[0,T]\times B_{R}}\|\sigma_{n}(t,x)-\sigma(t,x)\|^{2}$
$\quad+(4TC^{2}+4C_{\mathrm{BDG}}C^{2})\int_{0}^{T}\mathbb{E}\left[\sup_{u\in[0,s\wedge\tau_{R}]}|X_{u}^{n}-X_{u}|^{2}\right]ds.$
Let $A_{n,R}=4T^{2}\sup_{(t,x)\in[0,T]\times B_{R}}|b_{n}(t,x)-b(t,x)|^{2}+4C_{\mathrm{BDG}}T\sup_{(t,x)\in[0,T]\times B_{R}}\|\sigma_{n}(t,x)-\sigma(t,x)\|^{2}$ and $K=4TC^{2}+4C_{\mathrm{BDG}}C^{2}$. By Gronwall’s inequality
$\mathbb{E}\left[\sup_{t\in[0,\tau_{R}]}|X_{t}^{n}-X_{t}|^{2}\right]\leq A_{n,R}e^{KT}.$
$\mathbb{E}|X_{T}^{n}-X_{T}|^{2}=\mathbb{E}\big{[}|X_{T}^{n}-X_{T}|^{2}\mathbf{1}_{\{\tau_{R}=T\}}\big{]}+\mathbb{E}\big{[}|X_{T}^{n}-X_{T}|^{2}\mathbf{1}_{\{\tau_{R}<T\}}\big{]},$
For fixed $R$, $A_{n,R}\to 0$ as $n\to\infty$. By (iii), $\sup_{n}\mathbb{E}\sup_{t\leq T}\|X_{t}^{n}\|^{2}+\mathbb{E}\sup_{t\leq T}\|X_{t}\|^{2}<\infty$, hence $\mathbb{P}(\tau_{R}<T)\to 0$ as $R\to\infty$, which yields $\mathbb{E}[|X_{T}^{n}-X_{T}|^{2}]\to 0$. The Malliavin derivative satisfies
$D_{t}X_{T}^{n}=Y_{T}^{n}(Y_{t}^{n})^{-1}\sigma_{n}(t,X_{t}^{n})\mathbf{1}_{\{t\leq T\}},\quad D_{t}X_{T}=Y_{T}Y_{t}^{-1}\sigma(t,X_{t})\mathbf{1}_{\{t\leq T\}},$
where the first variation processes satisfy
$dY_{t}^{n}=\partial_{x}b_{n}(t,X_{t}^{n})Y_{t}^{n}\,dt+\sum_{l=1}^{d}\partial_{x}\sigma_{n,l}(t,X_{t}^{n})Y_{t}^{n}\,dB_{t}^{l},\quad Y_{0}^{n}=I_{m}.$
Since $\partial_{x}b_{n}\to\partial_{x}b$ and $\partial_{x}\sigma_{n}\to\partial_{x}\sigma$ uniformly with uniform bounds, we obtain
$\mathbb{E}\left[\sup_{t\in[0,T]}\|Y_{t}^{n}-Y_{t}\|^{2}\right]\to 0,\quad\mathbb{E}\left[\sup_{t\in[0,T]}\|(Y_{t}^{n})^{-1}-Y_{t}^{-1}\|^{2}\right]\to 0.$
Decomposing the difference
$D_{t}X_{T}^{n}-D_{t}X_{T}$ $=Y_{T}^{n}[(Y_{t}^{n})^{-1}\sigma_{n}(t,X_{t}^{n})-Y_{t}^{-1}\sigma(t,X_{t})]\mathbf{1}_{\{t\leq T\}}$
$\quad+(Y_{T}^{n}-Y_{T})Y_{t}^{-1}\sigma(t,X_{t})\mathbf{1}_{\{t\leq T\}}.$
Therefore
$|D_{t}X_{T}^{n}-D_{t}X_{T}|^{2}$ $\leq 2\|Y_{T}^{n}\|^{2}\|(Y_{t}^{n})^{-1}\sigma_{n}(t,X_{t}^{n})-Y_{t}^{-1}\sigma(t,X_{t})\|^{2}$
$\quad+2\|Y_{T}^{n}-Y_{T}\|^{2}\|Y_{t}^{-1}\sigma(t,X_{t})\|^{2}.$
Setting $I_{1}^{n}$ and $I_{2}^{n}$ as the integrals of these terms, and using Cauchy-Schwarz with $\sup_{n}\mathbb{E}[\|Y_{T}^{n}\|^{4}]<\infty$
$I_{2}^{n}\leq 2(\mathbb{E}[\|Y_{T}^{n}-Y_{T}\|^{4}])^{1/2}\left(\mathbb{E}\left[\left(\int_{0}^{T}\|Y_{t}^{-1}\sigma(t,X_{t})\|^{2}dt\right)^{2}\right]\right)^{1/2}\to 0.$
Similarly, $I_{1}^{n}\to 0$ since the integrand converges to zero and the first factor is bounded. Thus
$\mathbb{E}\left[\int_{0}^{T}|D_{t}X_{T}^{n}-D_{t}X_{T}|^{2}dt\right]\to 0.$
We have
$\gamma^{n}_{ij}-\gamma_{ij}=\int_{0}^{T}(D_{t}X_{T}^{n,i}D_{t}X_{T}^{n,j}-D_{t}X_{T}^{i}D_{t}X_{T}^{j})dt.$
Decomposing the product difference
$\gamma^{n}_{ij}-\gamma_{ij}=\int_{0}^{T}\left[(D_{t}X_{T}^{n,i}-D_{t}X_{T}^{i})D_{t}X_{T}^{n,j}+D_{t}X_{T}^{i}(D_{t}X_{T}^{n,j}-D_{t}X_{T}^{j})\right]dt,$
we obtain
$\mathbb{E}\|\gamma^{n}-\gamma\|^{2}\leq C\sum_{i,j}\bigg{(}$ $\mathbb{E}\left[\int_{0}^{T}|D_{t}X_{T}^{n,i}-D_{t}X_{T}^{i}|^{2}|D_{t}X_{T}^{n,j}|^{2}dt\right]$
$+\mathbb{E}\left[\int_{0}^{T}|D_{t}X_{T}^{n,j}-D_{t}X_{T}^{j}|^{2}|D_{t}X_{T}^{i}|^{2}dt\right]\bigg{)}.$
We had, $D_{t}X_{T}^{n}\to D_{t}X_{T}$ in $L^{2}([0,T]\times\Omega)$. Combined with the uniform $L^{4}$ bounds for $D_{t}X_{T}^{n}$ (which hold under the $C^{2}$ assumptions via $D_{t}X_{T}^{n}=Y_{T}^{n}(Y_{t}^{n})^{-1}\sigma_{n}(t,X_{t}^{n})$), the right-hand side converges to 0. Since $\gamma^{n},\gamma\geq\lambda I_{m}$, we have the operator norm bound
$\|(\gamma^{n})^{-1}-\gamma^{-1}\|\leq\lambda^{-2}\|\gamma^{n}-\gamma\|,$
which implies $\mathbb{E}[\|(\gamma^{n})^{-1}-\gamma^{-1}\|^{2}]\leq\lambda^{-4}\mathbb{E}[\|\gamma^{n}-\gamma\|^{2}]\to 0$. We decompose
$u_{k}^{n}(t)-u_{k}(t)$ $=\sum_{j=1}^{m}(\gamma^{n})^{-1}_{jk}(D_{t}X_{T}^{n,j}-D_{t}X_{T}^{j})$
$\quad+\sum_{j=1}^{m}[(\gamma^{n})^{-1}_{jk}-\gamma^{-1}_{jk}]D_{t}X_{T}^{j}.$
Squaring and using $\sum_{j}|(\gamma^{n})^{-1}_{jk}|^{2}\leq\|(\gamma^{n})^{-1}\|^{2}\leq\lambda^{-2}$
$|u_{k}^{n}(t)-u_{k}(t)|^{2}$ $\leq 2\lambda^{-2}\sum_{j}|D_{t}X_{T}^{n,j}-D_{t}X_{T}^{j}|^{2}$
$\quad+2\|(\gamma^{n})^{-1}-\gamma^{-1}\|^{2}\sum_{j}|D_{t}X_{T}^{j}|^{2}.$
The integrand converges to 0 in $L^{1}([0,T]\times\Omega)$ by Steps 2 and 3, and is dominated by an integrable function since $\mathrm{Tr}(\gamma)\in L^{2}(\Omega)$. By dominated convergence
$\mathbb{E}\left[\int_{0}^{T}|u_{k}^{n}(t)-u_{k}(t)|^{2}dt\right]\to 0.$
4 Variation processes
Given a stochastic process $X_{t}$ evolving from the initial condition $x$ in $\mathbb{R}^{m}$ we define the first variation process as $Y_{t}=\partial X_{t}/\partial x\in\mathbb{R}^{m\times m}$. As is well known, $Y_{t}$ represents the sensitivity of $X_{t}$ relative to small variations in the initial condition $x$. Differentiating the SDE (2.1) with respect to $x$ yields
$dY_{t}=\partial_{x}b(t,X_{t})Y_{t}\,dt+\sum_{l=1}^{d}\partial_{x}\sigma^{l}(t,X_{t})Y_{t}\,dB_{t}^{l},\quad Y_{0}=I_{m},$
where $\partial_{x}b(t,X_{t})\in\mathbb{R}^{m\times m}$, with $\left[\partial_{x}b(t,X_{t})\right]_{i,j}={\partial b^{i}(t,X_{t})}/{\partial x_{j}}$, $\partial_{x}\sigma^{l}(t,X_{t})\in\mathbb{R}^{m\times m}$, with $\left[\partial_{x}\sigma^{l}(t,X_{t})\right]_{i,j}={\sigma^{i,l}(t,X_{t})}/{\partial x_{j}}$, and $I_{m}$ is the $m\times m$ identity matrix. The Malliavin derivative $D_{t}X_{T}$ is the response of $X_{T}$ to a perturbation in the Brownian motion at time $t$. For $t\leq T$
$D_{t}X_{T}^{j}=\left[Y_{T}Y_{t}^{-1}\sigma(t,X_{t})\right]^{j}.$
For $t>T$, $D_{t}X_{T}=0$ (future perturbations don’t affect $X_{T}$). This follows because $D_{t}X_{s}=0$ for $s<t$, and $D_{t}X_{t}=\sigma(t,X_{t})$, with the perturbation propagating via $Y_{T,t}=Y_{T}Y_{t}^{-1}$.
The second variation process, denoted $Z_{t}=\partial^{2}X_{t}/\partial x^{2}$, is a third-order tensor in $\mathbb{R}^{m\times m\times m}$. It represents the second-order sensitivities of the state process $X_{t}$ with respect to the initial condition $x$. Each component of $Z_{t}$, written as $Z_{t}^{i,p,q}$, corresponds to the second partial derivative $\partial^{2}X_{t}^{i}/\partial x_{p}\partial x_{q}$, where $i,p,q=1,\ldots,m$. This process evolves according to the following SDE
$dZ_{t}$ $=\left[\partial_{xx}b(t,X_{t})(Y_{t}\otimes Y_{t})+\partial_{x}b(t,X_{t})Z_{t}\right]dt$
$\quad+\sum_{l=1}^{d}\left[\partial_{xx}\sigma^{l}(t,X_{t})(Y_{t}\otimes Y_{t})+\partial_{x}\sigma^{l}(t,X_{t})Z_{t}\right]dB_{t}^{l},$
With the initial condition $Z_{0}=0$, the stochastic differential equation for the second variation process $Z_{t}$ is characterised by a collection of terms that describe the system’s dynamics. The Hessian tensor of the drift coefficient $b$, denoted $\partial_{xx}b(t,X_{t})\in\mathbb{R}^{m\times m\times m}$, has components defined as $\left[\partial_{xx}b\right]_{i,j,k}=\partial^{2}b^{i}(t,X_{t})/(\partial x_{j}\partial x_{k})$, representing the second derivatives of the $i$-th component of $b$ with respect to the spatial variables $x_{j}$ and $x_{k}$. Likewise, the Hessian tensor of the $l$-th column of the diffusion coefficient $\sigma$, expressed as $\partial_{xx}\sigma^{l}(t,X_{t})\in\mathbb{R}^{m\times m\times m}$, accounts for higher-order effects in the stochastic terms. The first variation process $Y_{t}\in\mathbb{R}^{m\times m}$ captures the first-order sensitivities of the state process $X_{t}$ with respect to the initial condition $x$. This leads to the tensor product $Y_{t}\otimes Y_{t}$, a fourth-order tensor in $\mathbb{R}^{m\times m\times m\times m}$, which interacts with the Hessian tensors through contraction, enriching the structure of the second-variation SDE. The Jacobian matrices $\partial_{x}b(t,X_{t})\in\mathbb{R}^{m\times m}$ and $\partial_{x}\sigma^{l}(t,X_{t})\in\mathbb{R}^{m\times m}$ of $b$ and $\sigma^{l}$, respectively, provide the first-order spatial dependencies of the drift and diffusion coefficients. Driving the stochastic nature of the system, $B_{t}^{l}$ represents the $l$-th component of a $d$-dimensional Brownian motion, introducing randomness into the evolution of $Z_{t}$. To clarify the tensor contraction, the term $\partial_{xx}b(t,X_{t})(Y_{t}\otimes Y_{t})$ for each component $Z_{t}^{i,p,q}$ is
$\left[\partial_{xx}b(t,X_{t})(Y_{t}\otimes Y_{t})\right]^{i,p,q}=\sum_{j,k=1}^{m}\frac{\partial^{2}b^{i}(t,X_{t})}{\partial x_{j}\partial x_{k}}Y_{t}^{j,p}Y_{t}^{k,q}.$
Similarly, the term $\partial_{x}b(t,X_{t})Z_{t}$ is
$\left[\partial_{x}b(t,X_{t})Z_{t}\right]^{i,p,q}=\sum_{r=1}^{m}\frac{\partial b^{i}(t,X_{t})}{\partial x_{r}}Z_{t}^{r,p,q}.$
The diffusion terms follow the same structure with $\sigma^{l}$ replacing $b$. By the classical existence‐and‐uniqueness theory for Itô–SDEs, together with the fact that all first and second‐order derivatives of $b$ and $\sigma$ are uniformly bounded (so that the coefficient‐processes in the linear equation are globally Lipschitz and square‐integrable), one obtains a unique strong solution $Z_{t}$ (see, e.g., Kunita [10], Ch. 4).
5 Proof of Theorem 2.1
In this section we prove our main result, i.e., Theorem 2.1. The proof relies on the results we obtained in Section 3 and Section 4 for covering vector fields and variation processes to define the gradient of the log density for general stochastic differential equations using a Bismut-type formula. This formulation enables the computation of score functions for general nonlinear diffusion processes governed by stochastic differential equations. In this section, we first establish several preliminary results before proceeding to the proof of our main Theorem 2.1.
We begin by considering the representation of the Skorokhod integral $\delta(u_{k})$. To this end, we first refer to Theorem 3.2.9 in [15], which provides a decomposition of the Skorokhod integral of a random field composed with a random variable. We also refer to the seminal works [17, 18], which study generalised stochastic integrals and anticipating integrals in Malliavin calculus. The result arises in the context of substitution formulae for stochastic integrals. Consider a random field $u=\{u_{t}(x):0\leq t\leq T,x\in\mathbb{R}^{m}\}$ with $u_{t}(x)\in\mathbb{R}^{d}$, which is square integrable and adapted for each $x\in\mathbb{R}^{m}$. For each $x$, one can define the Itô integral $\int_{0}^{T}u_{t}(x)\cdot dB_{t}$. Given an $m$-dimensional random variable $F:\Omega\to\mathbb{R}^{m}$, Theorem 3.2.9 addresses the Skorokhod integrability of the nonadapted process $u(F)=\{u_{t}(F),0\leq t\leq T\}$ and provides a formula for its Skorokhod integral under the following conditions
  • (h1)
    For each $x\in\mathbb{R}^{m}$ and $t\in[0,T]$, $u_{t}(x)$ is $\mathcal{F}_{t}$-measurable.
  • (h2)
    There exist constants $p\geq 2$ and $\alpha>m$ such that
    $E(|u_{t}(x)-u_{t}(y)|^{p})\leq C_{t,K}|x-y|^{\alpha},$
    for all $|x|,|y|\leq K$, $K>0$, where $\int_{0}^{T}C_{t,K}dt<\infty$. Moreover,
    $\int_{0}^{T}E(|u_{t}(0)|^{2})dt<\infty.$
  • (h3)
    For each $(t,\omega)$, the mapping $x\mapsto u_{t}(x)$ is continuously differentiable, and for each $K>0$,
    $\int_{0}^{T}E\left(\sup_{|x|\leq K}|\nabla u_{t}(x)|^{q}\right)dt<\infty,$
    where $q\geq 4$ and $q>m$.
For the reader’s convenience, we recall Theorem 3.2.9 from [15] below.
Theorem 5.1 (Theorem 3.2.9, [15]).
For a random field $u=\{u_{t}(x):0\leq t\leq T,x\in\mathbb{R}^{m}\}$ with $u_{t}(x)\in\mathbb{R}^{d}$, and a random variable $F:\Omega\to\mathbb{R}^{m}$ such that $F^{i}\in\mathbb{D}^{1,4}_{\text{loc}}$ for $1\leq i\leq m$, assume $u$ satisfies the conditions (h1) and (h3) for Skorokhod integrability. Then, the composition $u(F)=\{u_{t}(F),0\leq t\leq T\}$ belongs to $(\text{Dom}(\delta))_{\text{loc}}$, and the Skorokhod integral of $u(F)$ is given by
$\delta(u(F))=\left.\int_{0}^{T}u_{t}(x)\cdot dB_{t}\right|_{x=F}-\sum_{j=1}^{m}\int_{0}^{T}\partial_{j}u_{t}(F)\cdot D_{t}F^{j}\,dt,$
where $B_{t}$ is a $d$-dimensional Brownian motion, $\partial_{j}u_{t}(x)=\partial u_{t}(x)/\partial x_{j}$ is the partial derivative of $u_{t}(x)$ with respect to the $j$-th component of $x$, $D_{t}F^{j}$ is the Malliavin derivative of the $j$-th component of $F$, and $\left.\int_{0}^{T}u_{t}(x)\cdot dB_{t}\right|_{x=F}$ denotes the Itô integral evaluated at $x=F$. We note that no smoothness in the sense of Malliavin calculus is required on the process $u_{t}(x)$ itself, but the above conditions ensure the integrability of $u(F)$ in the Skorokhod sense. Furthermore, the operator $\delta$ is not known to be local in $\text{Dom}(\delta)$, and thus the value of $\delta(u(F))$ may depend on the particular localising sequence used in the definition of $(\text{Dom}(\delta))_{\text{loc}}$.
5.1 Useful lemmas
In this section, we state and prove some key results that are useful for deriving the score function formula (2.3)-(2.4) for nonlinear SDEs in terms of the first and second variation processes.
Lemma 5.2 (SDE for the inverse first variation process).
Let $Y_{t}$ be the first variation process. The inverse $Y_{t}^{-1}$ satisfies the SDE
$dY_{t}^{-1}$ $=-Y_{t}^{-1}\partial_{x}b(t,X_{t})\,dt-\sum_{l=1}^{d}Y_{t}^{-1}\partial_{x}\sigma^{l}(t,X_{t})\,dB_{t}^{l}+\sum_{l=1}^{d}Y_{t}^{-1}\left(\partial_{x}\sigma^{l}(t,X_{t})\right)^{2}\,dt,$
with initial condition $Y_{0}^{-1}=I_{m}$, where $\left(\partial_{x}\sigma^{l}(t,X_{t})\right)^{2}=\partial_{x}\sigma^{l}(t,X_{t})\partial_{x}\sigma^{l}(t,X_{t})$.
Proof.
Since $Y_{t}Y_{t}^{-1}=I_{m}$ is constant, its differential is zero, i.e.,
$d(Y_{t}Y_{t}^{-1})$ $=dY_{t}Y_{t}^{-1}+Y_{t}dY_{t}^{-1}+d[Y_{t},Y_{t}^{-1}]=0.$
Given the first-variation SDE
$dY_{t}$ $=\partial_{x}b(t,X_{t})Y_{t}\,dt+\sum_{l=1}^{d}\partial_{x}\sigma^{l}(t,X_{t})Y_{t}\,dB_{t}^{l},$ (5.1)
we compute
$dY_{t}Y_{t}^{-1}$ $=\left(\partial_{x}b(t,X_{t})Y_{t}\,dt+\sum_{l=1}^{d}\partial_{x}\sigma^{l}(t,X_{t})Y_{t}\,dB_{t}^{l}\right)Y_{t}^{-1}$
$=\partial_{x}b(t,X_{t})\,dt+\sum_{l=1}^{d}\partial_{x}\sigma^{l}(t,X_{t})\,dB_{t}^{l},$
where we used $Y_{t}Y_{t}^{-1}=I_{m}$. We assume that
$dY_{t}^{-1}$ $=\mu_{t}\,dt+\sum_{l=1}^{d}\nu_{t}^{l}\,dB_{t}^{l}.$
We need to show $Y_{t}^{-1}$ is an Itô process. To this end, consider the SDE (5.1) with $Y_{0}=I_{m}$. The coefficients $\partial_{x}b(t,X_{t})$ and $\partial_{x}\sigma^{l}(t,X_{t})$ are assumed to be bounded and measurable. The SDE (5.1) is linear, and under standard conditions (e.g., Lipschitz continuity of the coefficients), the solution $Y_{t}$ exists, is unique, and is a semimartingale — specifically, an Itô process — adapted to the filtration generated by the Brownian motions $B_{t}^{l}$. Moreover, since $Y_{0}=I_{m}$ is invertible and the coefficients satisfy regularity conditions, $Y_{t}$ remains invertible almost surely for all $t\geq 0$. Define the function $f:\text{GL}(m,\mathbb{R})\to\text{GL}(m,\mathbb{R})$ by $f(A)=A^{-1}$, where $\text{GL}(m,\mathbb{R})$ is the group of $m\times m$ invertible matrices. The map $f$ is smooth (infinitely differentiable) on $\text{GL}(m,\mathbb{R})$, with first derivative $Df(A)H=-A^{-1}HA^{-1}$ and second derivative terms involving higher-order products. Applying Itô’ s formula to $Y_{t}^{-1}=f(Y_{t})$, where $Y_{t}$ is an Itô process, yields a stochastic differential of the form
$dY_{t}^{-1}$ $=\mu_{t}\,dt+\sum_{l=1}^{d}\nu_{t}^{l}\,dB_{t}^{l},$
where $\mu_{t}$ and $\nu_{t}^{l}$ are adapted processes derived from the drift and diffusion terms of $Y_{t}$. Thus, $Y_{t}^{-1}$ is itself an Itô process, and we proceed to determine $\mu_{t}$ and $\nu_{t}^{l}$. Continuing with the proof
$Y_{t}dY_{t}^{-1}$ $=Y_{t}\mu_{t}\,dt+\sum_{l=1}^{d}Y_{t}\nu_{t}^{l}\,dB_{t}^{l}.$
The quadratic covariation is
$d[Y_{t},Y_{t}^{-1}]$ $=\sum_{l=1}^{d}\left(\partial_{x}\sigma^{l}(t,X_{t})Y_{t}\right)\nu_{t}^{l}\,dt.$
Substituting into the differential
$d(Y_{t}Y_{t}^{-1})$ $=\left(\partial_{x}b(t,X_{t})+Y_{t}\mu_{t}+\sum_{l=1}^{d}\partial_{x}\sigma^{l}(t,X_{t})Y_{t}\nu_{t}^{l}\right)dt$
$\quad+\sum_{l=1}^{d}\left(\partial_{x}\sigma^{l}(t,X_{t})+Y_{t}\nu_{t}^{l}\right)dB_{t}^{l}$
$=0,$
and equating the coefficients yields
$dB_{t}^{l}\text{-term}:Y_{t}\nu_{t}^{l}+\partial_{x}\sigma^{l}(t,X_{t})=0,\quad\Rightarrow\quad\nu_{t}^{l}=-Y_{t}^{-1}\partial_{x}\sigma^{l}(t,X_{t}),$
$dt\text{-term}:\quad Y_{t}\mu_{t}+\partial_{x}b(t,X_{t})+\sum_{l=1}^{d}\partial_{x}\sigma^{l}(t,X_{t})Y_{t}(-Y_{t}^{-1}\partial_{x}\sigma^{l}(t,X_{t}))=0.$
Simplifying
$\sum_{l=1}^{d}\partial_{x}\sigma^{l}(t,X_{t})Y_{t}(-Y_{t}^{-1}\partial_{x}\sigma^{l}(t,X_{t}))$ $=-\sum_{l=1}^{d}\partial_{x}\sigma^{l}(t,X_{t})\partial_{x}\sigma^{l}(t,X_{t}),$
$Y_{t}\mu_{t}+\partial_{x}b(t,X_{t})-\sum_{l=1}^{d}\partial_{x}\sigma^{l}(t,X_{t})\partial_{x}\sigma^{l}(t,X_{t})$ $=0,$
$\mu_{t}$ $=Y_{t}^{-1}\left(-\partial_{x}b(t,X_{t})+\sum_{l=1}^{d}\left(\partial_{x}\sigma^{l}(t,X_{t})\right)^{2}\right).$
Thus
$dY_{t}^{-1}$ $=Y_{t}^{-1}\left(-\partial_{x}b(t,X_{t})+\sum_{l=1}^{d}\left(\partial_{x}\sigma^{l}(t,X_{t})\right)^{2}\right)dt-\sum_{l=1}^{d}Y_{t}^{-1}\partial_{x}\sigma^{l}(t,X_{t})\,dB_{t}^{l}.$
The initial condition $Y_{0}^{-1}=I_{m}$ holds since $Y_{0}=I_{m}$.
Lemma 5.3 (Malliavin derivative of inverse matrices).
Let $A$ be a random $m\times m$ matrix that is invertible almost surely, and assume there exists $\lambda>0$ such that $A\geq\lambda I_{m}$ almost surely (uniform ellipticity). Suppose further that $A$ and $A^{-1}$ are Malliavin differentiable. Then, for each $t\in[0,T]$,
$D_{t}(A^{-1})=-A^{-1}(D_{t}A)A^{-1}.$
Proof.
Since $AA^{-1}=I$, applying the Malliavin derivative yields
$D_{t}(AA^{-1})=(D_{t}A)A^{-1}+A(D_{t}A^{-1})=D_{t}I=0.$
Consequently,
$A(D_{t}A^{-1})=-(D_{t}A)A^{-1}.$
Multiplying on the left by $A^{-1}$ gives the desired result.
Remark 5.4 (The role of integrability and ellipticity conditions).
While Lemma 5.3 is purely algebraic, ensuring that $A^{-1}$ (and the right-hand side) belongs to the Malliavin–Sobolev space $\mathbb{D}^{1,2}$ requires controlling the integrability of $\|A^{-1}\|$. If $A\geq\lambda I_{m}$ almost surely for some deterministic $\lambda>0$, then $\|A^{-1}\|\leq\lambda^{-1}$ and all polynomial moments of $A^{-1}$ are finite. Under this bound, every entry of $A^{-1}$ lies in $\mathbb{D}^{\infty}$, and the componentwise identity
$D(A^{-1})^{ij}=-\sum_{k,\ell=1}^{m}(A^{-1})^{ik}(A^{-1})^{\ell j}\,DA^{k\ell},$
established in [15, Lemma 2.1.6], is justified. Without such a lower bound on the eigenvalues, the inverse may fail to be integrable. In dimension one, for instance, $A=\int_{0}^{T}B_{s}^{2}\,ds>0$ satisfies $A^{-q}\in L^{1}(\Omega)$ only for $q<\frac{1}{2}$; thus $A^{-1}\notin\mathbb{D}^{1,2}$ and Lemma 2.1.6 cannot be applied. Uniform ellipticity (or, more generally, the moment hypothesis $|\det A|^{-1}\in L^{p}$ for all $p\geq 1$ required in [15, Lemma 2.1.6]) prevents this pathology.
Lemma 5.5 (Commutativity of Malliavin and partial derivatives).
Let $X_{t}=X_{t}(x)$ be the solution to the stochastic differential equation
$\mathrm{d}X_{t}=b(t,X_{t})\,\mathrm{d}t+\sigma(t,X_{t})\,\mathrm{d}B_{t},\quad X_{0}=x,\quad 0\leq t\leq T,$
where $B_{t}$ is a $d$-dimensional standard Brownian motion and $x\in\mathbb{R}^{m}$. Assume the coefficients satisfy $b,\sigma\in\mathcal{C}^{2}_{\!b}([0,T]\times\mathbb{R}^{m})$ (all partial derivatives up to second order are continuous and bounded). Let $Y_{t}=\partial X_{t}/\partial x$ denote the first variation process. Then, for every $0\leq t\leq T$, where $D_{t}$ denotes the Malliavin derivative,
$D_{t}\left(\frac{\partial X_{T}}{\partial x}\right)=\frac{\partial}{\partial x}(D_{t}X_{T}).$
Proof.
We compute both sides explicitly to verify equality. The SDE for $X_{t}$ can be rewritten in the integral form
$X_{t}$ $=x+\int_{0}^{t}b(s,X_{s})\,ds+\int_{0}^{t}\sigma(s,X_{s})\,dB_{s}.$
The first variation process $Y_{t}=\partial X_{t}/\partial x$ satisfies
$dY_{t}$ $=\partial_{x}b(t,X_{t})Y_{t}\,dt+\sum_{l=1}^{d}\partial_{x}\sigma^{l}(t,X_{t})Y_{t}\,dB_{t}^{l},\quad Y_{0}=I_{m}.$
We have seen that the inverse $Y_{t}^{-1}$ satisfies the SDE
$dY_{t}^{-1}$ $=-Y_{t}^{-1}\partial_{x}b(t,X_{t})\,dt$
$\quad-\sum_{l=1}^{d}Y_{t}^{-1}\partial_{x}\sigma^{l}(t,X_{t})\,dB_{t}^{l}$
$\quad+\sum_{l=1}^{d}Y_{t}^{-1}\left(\partial_{x}\sigma^{l}(t,X_{t})\right)^{2}\,dt,\quad Y_{0}^{-1}=I_{m}$
while the second variation process satisfies
$dZ_{t}$ $=\left[\partial_{xx}b(t,X_{t})(Y_{t}\otimes Y_{t})+\partial_{x}b(t,X_{t})Z_{t}\right]dt$
$\quad+\sum_{l=1}^{d}\left[\partial_{xx}\sigma^{l}(t,X_{t})(Y_{t}\otimes Y_{t})+\partial_{x}\sigma^{l}(t,X_{t})Z_{t}\right]dB_{t}^{l},\quad Z_{0}=0.$
$D_{t}X_{T}$ $=Y_{T}Y_{t}^{-1}\sigma(t,X_{t}).$
Then
$\frac{\partial}{\partial x}(D_{t}X_{T})$ $=\frac{\partial}{\partial x}\left(Y_{T}Y_{t}^{-1}\sigma(t,X_{t})\right)$
$=\frac{\partial Y_{T}}{\partial x}\left(Y_{t}^{-1}\sigma(t,X_{t})\right)+Y_{T}\frac{\partial}{\partial x}\left(Y_{t}^{-1}\sigma(t,X_{t})\right).$
Since $\partial Y_{T}/\partial x=Z_{T}$ we have
$\frac{\partial Y_{T}}{\partial x}\left(Y_{t}^{-1}\sigma(t,X_{t})\right)$ $=Z_{T}Y_{t}^{-1}\sigma(t,X_{t}).$
For the second term
$\frac{\partial}{\partial x}\left(Y_{t}^{-1}\sigma(t,X_{t})\right)$ $=\frac{\partial Y_{t}^{-1}}{\partial x}\sigma(t,X_{t})+Y_{t}^{-1}\frac{\partial\sigma(t,X_{t})}{\partial x},$
where $\partial Y_{t}^{-1}/\partial x=-Y_{t}^{-1}Z_{t}Y_{t}^{-1}$ and $\partial\sigma(t,X_{t})/\partial x=\partial_{x}\sigma(t,X_{t})Y_{t}$. Thus
$\frac{\partial}{\partial x}\left(Y_{t}^{-1}\sigma(t,X_{t})\right)$ $=-\left(Y_{t}^{-1}Z_{t}Y_{t}^{-1}\right)\sigma(t,X_{t})+Y_{t}^{-1}\left(\partial_{x}\sigma(t,X_{t})Y_{t}\right),$
$\frac{\partial}{\partial x}(D_{t}X_{T})$ $=Z_{T}Y_{t}^{-1}\sigma(t,X_{t})+Y_{T}\left(-\left(Y_{t}^{-1}Z_{t}Y_{t}^{-1}\right)\sigma(t,X_{t})+Y_{t}^{-1}\left(\partial_{x}\sigma(t,X_{t})Y_{t}\right)\right).$
Moreover, since $\partial X_{T}/\partial x=Y_{T}$ we have
$Y_{T}$ $=I_{m}+\int_{0}^{T}\partial_{x}b(s,X_{s})Y_{s}\,ds+\int_{0}^{T}\partial_{x}\sigma(s,X_{s})Y_{s}\,dB_{s},$
$D_{t}Y_{T}$ $=\partial_{x}\sigma(t,X_{t})Y_{t}+\int_{t}^{T}\left[\partial_{x}b(s,X_{s})D_{t}Y_{s}+\partial_{xx}b(s,X_{s})\left(Y_{s}Y_{t}^{-1}\sigma(t,X_{t})\right)Y_{s}\right]ds$
$\quad+\int_{t}^{T}\left[\partial_{x}\sigma(s,X_{s})D_{t}Y_{s}+\partial_{xx}\sigma(s,X_{s})\left(Y_{s}Y_{t}^{-1}\sigma(t,X_{t})\right)Y_{s}\right]dB_{s}.$
The solution is
$D_{t}Y_{T}$ $=Z_{T}Y_{t}^{-1}\sigma(t,X_{t})-Y_{T}Y_{t}^{-1}Z_{t}Y_{t}^{-1}\sigma(t,X_{t})+Y_{T}Y_{t}^{-1}\partial_{x}\sigma(t,X_{t})Y_{t}.$
Clearly the expressions are equal, confirming that
$D_{t}\left(\frac{\partial X_{T}}{\partial x}\right)$ $=\frac{\partial}{\partial x}\left(D_{t}X_{T}\right).$
The regularity conditions ensure all derivatives and integrals are well-defined.
Lemma 5.6 (Malliavin derivative for the first variation process).
For $t\leq T$, the Malliavin derivative of the first variation process $Y_{T}$ is given by
$D_{t}Y_{T}=$ $\,Z_{T}\,Y_{t}^{-1}\,\sigma(t,X_{t})$
$-\,Y_{T}\,Y_{t}^{-1}\,Z_{t}\,Y_{t}^{-1}\,\sigma(t,X_{t})$
$+\,Y_{T}\,Y_{t}^{-1}\,\partial_{x}\sigma(t,X_{t})\,Y_{t}.$
Proof.
Consider the SDE satisfied by the first variation process
$dY_{s}=\partial_{x}b(s,X_{s})\,Y_{s}\,ds\;+\;\partial_{x}\sigma(s,X_{s})\,Y_{s}\,dB_{s},\qquad Y_{0}=I.$
Recall that $Y_{s}=\partial X_{s}/\partial x$. The Malliavin derivative $D_{t}Y_{T}$ represents the sensitivity of $Y_{T}$ to a perturbation in the Brownian motion at time $t$. Since $Y_{T}=\partial X_{T}/\partial x$ we have
$D_{t}Y_{T}$ $=D_{t}\left(\frac{\partial X_{T}}{\partial x}\right)=\frac{\partial}{\partial x}\left(D_{t}X_{T}\right).$
This identity follows from interchanging the partial derivative w.r.t. $x$ and the Malliavin derivative $D_{t}$. We use the known expression $D_{t}X_{T}=Y_{T}\,\bigl{(}Y_{t}^{-1}\,\sigma(t,X_{t})\bigr{)}$ which is valid for all $t\leq T$. Define $W_{t}=Y_{t}^{-1}\,\sigma(t,X_{t})$. Thus,
$D_{t}X_{T}$ $=Y_{T}\,W_{t}.$
Differentiate this equation with respect to the initial condition $x$ to obtain
$\frac{\partial}{\partial x}(D_{t}X_{T})=\frac{\partial Y_{T}}{\partial x}W_{t}+\ Y_{T}\,\frac{\partial W_{t}}{\partial x}.$
By definition of second variation process $\partial Y_{T}/\partial x=Z_{T}$. Hence
$D_{t}Y_{T}=Z_{T}\,W_{t}+Y_{T}\frac{\partial W_{t}}{\partial x}.$ (5.2)
Recalling that $W_{t}=Y_{t}^{-1}\,\sigma(t,X_{t})$ we can write
$\frac{\partial W_{t}}{\partial x}$ $=\frac{\partial}{\partial x}\left(Y_{t}^{-1}\,\sigma(t,X_{t})\right)$
$=\left(\frac{\partial Y_{t}^{-1}}{\partial x}\right)\,\sigma(t,X_{t})+Y_{t}^{-1}\,\frac{\partial\sigma(t,X_{t})}{\partial x}.$
Each of these two terms comes from the product rule (now for partial derivatives w.r.t. $x$). Since $Y_{t}\,Y_{t}^{-1}=I$, differentiating both sides w.r.t. $x$ yields $\partial Y_{t}/\partial x\,Y_{t}^{-1}+Y_{t}\,\partial Y_{t}^{-1}/\partial x=0.$ Hence
$\frac{\partial Y_{t}^{-1}}{\partial x}$ $=-Y_{t}^{-1}\,\left(\frac{\partial Y_{t}}{\partial x}\right)\,Y_{t}^{-1}=-Y_{t}^{-1}Z_{t}Y_{t}^{-1}.$
We also have $\partial X_{t}/\partial x=Y_{t}$. Thus by chain rule
$\frac{\partial\sigma(t,X_{t})}{\partial x}$ $=\partial_{x}\sigma(t,X_{t})\,Y_{t}.$
Therefore,
$\frac{\partial W_{t}}{\partial x}$ $=-Y_{t}^{-1}Z_{t}Y_{t}^{-1}\sigma(t,X_{t})+Y_{t}^{-1}\partial_{x}\sigma(t,X_{t})Y_{t}.$
This completes the computation of $\partial W_{t}/\partial x$. Putting this result back into (5.2) yields
$D_{t}Y_{T}$ $=Z_{T}\,Y_{t}^{-1}\,\sigma(t,X_{t})\;+\;Y_{T}\Bigl{[}-\,Y_{t}^{-1}\,Z_{t}\,Y_{t}^{-1}\,\sigma(t,X_{t})\;+\;Y_{t}^{-1}\,\partial_{x}\sigma(t,X_{t})\,Y_{t}\Bigr{]}.$
Factor out common terms to rewrite it in the stated form
$D_{t}Y_{T}=$ $\,Z_{T}\,Y_{t}^{-1}\,\sigma(t,X_{t})-\,Y_{T}\,Y_{t}^{-1}\,Z_{t}\,Y_{t}^{-1}\,\sigma(t,X_{t})+\,Y_{T}\,Y_{t}^{-1}\,\partial_{x}\sigma(t,X_{t})\,Y_{t}.$
The second term $Y_{T}\,Y_{t}^{-1}\,Z_{t}\,Y_{t}^{-1}\,\sigma(t,X_{t})$ accounts for the appearance of the second variation $Z_{t}$ inside the inverse, while the final term $Y_{T}\,Y_{t}^{-1}\,\partial_{x}\sigma(t,X_{t})\,Y_{t}$ encodes the effect of differentiating $\sigma$ itself w.r.t. $x$.
Lemma 5.7 (Malliavin derivative formula for the inverse first variation process).
For the inverse first variation process $Y_{s}^{-1}$, the Malliavin derivative is given by
  • For $t\leq s$:
    $D_{t}Y_{s}^{-1}$ $=-Y_{s}^{-1}\left[Z_{s}Y_{t}^{-1}\sigma(t,X_{t})-Y_{s}Y_{t}^{-1}Z_{t}Y_{t}^{-1}\sigma(t,X_{t})+Y_{s}Y_{t}^{-1}\partial_{x}\sigma(t,X_{t})Y_{t}\right]Y_{s}^{-1}$
  • For $t>s$:
    $D_{t}Y_{s}^{-1}=0$
where $Z_{s}=\partial^{2}X_{s}/\partial x^{2}$ is the second variation process.
Proof.
We derive $D_{t}Y_{s}^{-1}$ by applying the Malliavin derivative to the identity $Y_{s}Y_{s}^{-1}=I_{m}$ and using the product rule. The proof splits into two cases based on the relationship between $t$ and $s$, and we leverage the expression for $D_{t}Y_{s}$ derived similarly to $D_{t}Y_{T}$. Let us begin with the case $t\leq s$. Since $Y_{s}Y_{s}^{-1}=I_{m}$ (the $m\times m$ identity matrix), we apply the Malliavin derivative $D_{t}$ to both sides
$D_{t}(Y_{s}Y_{s}^{-1})$ $=D_{t}(I_{m})=0.$
Using the product rule for Malliavin derivatives
$D_{t}(Y_{s}Y_{s}^{-1})=(D_{t}Y_{s})Y_{s}^{-1}+Y_{s}(D_{t}Y_{s}^{-1})=0.$
Rearranging to isolate $D_{t}Y_{s}^{-1}$
$Y_{s}(D_{t}Y_{s}^{-1})$ $=-(D_{t}Y_{s})Y_{s}^{-1},$
$D_{t}Y_{s}^{-1}$ $=-Y_{s}^{-1}(D_{t}Y_{s})Y_{s}^{-1}.$
To proceed, we need $D_{t}Y_{s}$. Since $Y_{s}=\partial X_{s}/\partial x$ and $t\leq s$, we adapt the derivation from the previous Lemma 5.6 for $D_{t}Y_{T}$, adjusting the endpoint from $T$ to $s$
$dY_{u}=\partial_{x}b(u,X_{u})Y_{u}\,du+\partial_{x}\sigma(u,X_{u})Y_{u}\,dB_{u},\quad 0\leq u\leq s,\qquad Y_{0}=I.$
We have
$D_{t}Y_{s}$ $=D_{t}\left(\frac{\partial X_{s}}{\partial x}\right)=\frac{\partial}{\partial x}(D_{t}X_{s}).$
For $t\leq s$, the Malliavin derivative of $X_{s}$ is
$D_{t}X_{s}$ $=Y_{s}Y_{t}^{-1}\sigma(t,X_{t}).$
Define $W_{t}=Y_{t}^{-1}\sigma(t,X_{t})$, so
$D_{t}X_{s}$ $=Y_{s}W_{t}.$
$\frac{\partial}{\partial x}(D_{t}X_{s})$ $=\frac{\partial}{\partial x}(Y_{s}W_{t})=\frac{\partial Y_{s}}{\partial x}W_{t}+Y_{s}\frac{\partial W_{t}}{\partial x}.$
Since $\partial Y_{s}/\partial x=Z_{s}$ (second variation process)
$D_{t}Y_{s}$ $=Z_{s}W_{t}+Y_{s}\frac{\partial W_{t}}{\partial x}.$
Recalling that $W_{t}=Y_{t}^{-1}\sigma(t,X_{t})$,
$\frac{\partial W_{t}}{\partial x}$ $=\frac{\partial Y_{t}^{-1}}{\partial x}\sigma(t,X_{t})+Y_{t}^{-1}\frac{\partial\sigma(t,X_{t})}{\partial x}.$
For $\partial Y_{t}^{-1}/\partial x$ differentiate $Y_{t}Y_{t}^{-1}=I$ to obtain
$\frac{\partial Y_{t}^{-1}}{\partial x}$ $=-Y_{t}^{-1}\frac{\partial Y_{t}}{\partial x}Y_{t}^{-1}=-Y_{t}^{-1}Z_{t}Y_{t}^{-1}.$
For $\partial\sigma(t,X_{t})/\partial x$, since $\partial X_{t}/\partial x=Y_{t}$, we have
$\frac{\partial\sigma(t,X_{t})}{\partial x}$ $=\partial_{x}\sigma(t,X_{t})Y_{t}.$
Therefore
$\frac{\partial W_{t}}{\partial x}$ $=-Y_{t}^{-1}Z_{t}Y_{t}^{-1}\sigma(t,X_{t})+Y_{t}^{-1}\partial_{x}\sigma(t,X_{t})Y_{t}$
and
$D_{t}Y_{s}$ $=Z_{s}Y_{t}^{-1}\sigma(t,X_{t})+Y_{s}\left(-Y_{t}^{-1}Z_{t}Y_{t}^{-1}\sigma(t,X_{t})+Y_{t}^{-1}\partial_{x}\sigma(t,X_{t})Y_{t}\right),$
$=Z_{s}Y_{t}^{-1}\sigma(t,X_{t})-Y_{s}Y_{t}^{-1}Z_{t}Y_{t}^{-1}\sigma(t,X_{t})+Y_{s}Y_{t}^{-1}\partial_{x}\sigma(t,X_{t})Y_{t}.$
Finally, substitute $D_{t}Y_{s}$ into $D_{t}Y_{s}^{-1}$ to obtain
$D_{t}Y_{s}^{-1}$ $=-Y_{s}^{-1}\left[Z_{s}Y_{t}^{-1}\sigma(t,X_{t})-Y_{s}Y_{t}^{-1}Z_{t}Y_{t}^{-1}\sigma(t,X_{t})+Y_{s}Y_{t}^{-1}\partial_{x}\sigma(t,X_{t})Y_{t}\right]Y_{s}^{-1}.$
Since $Y_{s}^{-1}$ is adapted to the filtration up to time $s$, and $t>s$, a perturbation in the Brownian motion at time $t$ does not affect $Y_{s}^{-1}$ (which depends only on information up to $s$). Thus
$D_{t}Y_{s}^{-1}$ $=0.$
This completes the proof, with the expression for $t\leq s$ matching the Lemma statement, and the zero result for $t>s$ reflecting the causality of the stochastic process.
5.2 Proof of Theorem 2.1
Having established all necessary preliminary results, we now proceed to prove our Theorem 2.1.
Proof.
Applying Theorem 5.1, the Skorokhod integral $\delta(u_{k})=\delta(u(F_{k}))$ is
$\delta(u_{k})=\left.\int_{0}^{T}u_{t}(x)\cdot dB_{t}\right|_{x=F_{k}}-\sum_{j=1}^{m}\int_{0}^{T}\partial_{j}u_{t}(F_{k})\cdot D_{t}F_{k}^{j}\,dt.$ (5.3)
This expression combines an Itô integral evaluated at $x=F_{k}$ with a correction term involving the partial derivatives of $u_{t}(x)$ evaluated at $x=F_{k}$ and the Malliavin derivatives of $F_{k}^{j}$. The first term in the expression for $\delta(u(F_{k}))$ is the Itô integral evaluated at $x=F_{k}$
$\left.\int_{0}^{T}u_{t}(x)\cdot dB_{t}\right|_{x=F_{k}},$
where for each fixed $x$, $u_{t}(x)=x^{\top}Y_{t}^{-1}\sigma(t,X_{t})$ is an adapted process, so $\int_{0}^{T}u_{t}(x)\cdot dB_{t}$ is a well-defined Itô integral, and after computing this integral, we evaluate it at $x=F_{k}=Y_{T}^{\top}\gamma_{X_{T}}^{-1}e_{k}$, which is $\mathcal{F}_{T}$-measurable. This term is computationally manageable because the integration is performed with respect to an adapted integrand for fixed $x$, and the randomness of $F_{k}$ is introduced only after the integration. With the redefined random field
$u_{t}(x)$ $=x^{\top}Y_{t}^{-1}\sigma(t,X_{t})=\sum_{i=1}^{m}x_{i}\left[Y_{t}^{-1}\sigma(t,X_{t})\right]_{i},$
the partial derivative with respect to $x_{j}$ is
$\partial_{j}u_{t}(x)$ $=\frac{\partial}{\partial x_{j}}u_{t}(x)=\left[Y_{t}^{-1}\sigma(t,X_{t})\right]_{j},$
since only the term involving $x_{j}$ depends on $x_{j}$. Therefore, evaluating at $x=F_{k}$
$\partial_{j}u_{t}(F_{k})$ $=\left[Y_{t}^{-1}\sigma(t,X_{t})\right]_{j}.$
This term will appear in the correction term of the Skorokhod integral decomposition. Before proceeding further, we state a general result from Lemma 5.3 for the Malliavin derivative of the inverse of a random matrix. To compute the Malliavin derivative $D_{t}F_{k}^{j}$, consider the new definition
$F_{k}=Y_{T}^{\top}\gamma_{X_{T}}^{-1}e_{k},\qquad F_{k}^{j}=e_{j}^{\top}Y_{T}^{\top}\gamma_{X_{T}}^{-1}e_{k},$
where $Y_{T}$ is the first variation process at time $T$, $\gamma_{X_{T}}$ is the Malliavin covariance matrix, and $e_{j},e_{k}$ are standard basis vectors. Applying the Malliavin derivative
$D_{t}F_{k}^{j}$ $=e_{j}^{\top}D_{t}(Y_{T}^{\top}\gamma_{X_{T}}^{-1})e_{k}.$
Using the product rule, we obtain
$D_{t}(Y_{T}^{\top}\gamma_{X_{T}}^{-1})$ $=(D_{t}Y_{T}^{\top})\gamma_{X_{T}}^{-1}+Y_{T}^{\top}D_{t}(\gamma_{X_{T}}^{-1}),$
$D_{t}F_{k}^{j}$ $=e_{j}^{\top}(D_{t}Y_{T}^{\top})\gamma_{X_{T}}^{-1}e_{k}+e_{j}^{\top}Y_{T}^{\top}D_{t}(\gamma_{X_{T}}^{-1})e_{k}.$
For $t\leq T$, the Malliavin derivative $D_{t}Y_{T}$ is given in Lemma 5.6
$D_{t}Y_{T}$ $=Z_{T}Y_{t}^{-1}\sigma(t,X_{t})-Y_{T}Y_{t}^{-1}Z_{t}Y_{t}^{-1}\sigma(t,X_{t})+Y_{T}Y_{t}^{-1}\partial_{x}\sigma(t,X_{t})Y_{t},$
where $Z_{t}$ is the second variation process. Taking the transpose
$D_{t}Y_{T}^{\top}$ $=\left[Z_{T}Y_{t}^{-1}\sigma(t,X_{t})-Y_{T}Y_{t}^{-1}Z_{t}Y_{t}^{-1}\sigma(t,X_{t})+Y_{T}Y_{t}^{-1}\partial_{x}\sigma(t,X_{t})Y_{t}\right]^{\top}.$
From Lemma 5.3 we have that for an invertible random matrix AA
$D_{t}(A^{-1})$ $=-A^{-1}(D_{t}A)A^{-1}.$
For the specific case of $A=\gamma_{X_{T}}$ this yields
$D_{t}(\gamma_{X_{T}}^{-1})$ $=-\gamma_{X_{T}}^{-1}(D_{t}\gamma_{X_{T}})\gamma_{X_{T}}^{-1},$
where
$\gamma_{X_{T}}$ $=\int_{0}^{T}Y_{T}Y_{s}^{-1}\sigma(s,X_{s})\sigma(s,X_{s})^{\top}(Y_{s}^{-1})^{\top}Y_{T}^{\top}\,ds,$
and $D_{t}\gamma_{X_{T}}$ requires computing the Malliavin derivative of the integrand. Thus,
$D_{t}F_{k}^{j}$ $=e_{j}^{\top}(D_{t}Y_{T}^{\top})\gamma_{X_{T}}^{-1}e_{k}+e_{j}^{\top}Y_{T}^{\top}D_{t}(\gamma_{X_{T}}^{-1})e_{k},$
$D_{t}F_{k}^{j}$ $=e_{j}^{\top}(D_{t}Y_{T}^{\top})\gamma_{X_{T}}^{-1}e_{k}-e_{j}^{\top}Y_{T}^{\top}\gamma_{X_{T}}^{-1}(D_{t}\gamma_{X_{T}})\gamma_{X_{T}}^{-1}e_{k},$
with $D_{t}Y_{T}^{\top}$ and $D_{t}\gamma_{X_{T}}$ as derived. We now proceed by computing the Malliavin derivative $D_{t}\Bigl{[}(Y_{T}\,Y_{s}^{-1}\,\sigma(s,X_{s}))^{p}\Bigr{]}$. To this end, we first recall Lemma 5.6 which gives the precise form of $D_{t}Y_{T}$. Afterwards, we use the product rule for Malliavin derivatives on the product $Y_{T}\,Y_{s}^{-1}\,\sigma(s,X_{s})$, distinguishing between the cases $t\leq s$ and $t>s$. Finally, we assemble these pieces to obtain the expression for $D_{t}(Y_{T}Y_{s}^{-1}\sigma(s,X_{s}))^{p}$. We provide reasoning for each step to clarify why each term appears and how the partial derivatives interact with the inverse processes. Let
$W_{s}^{p}\;=\;\Bigl{(}Y_{T}\,Y_{s}^{-1}\,\sigma(s,X_{s})\Bigr{)}^{p},$
i.e. the $p$-th component of the vector $Y_{T}\,Y_{s}^{-1}\,\sigma(s,X_{s})$. We want to find $D_{t}\bigl{(}W_{s}^{p}\bigr{)}$. Since
$W_{s}^{p}=\bigl{(}Y_{T}\,Y_{s}^{-1}\,\sigma(s,X_{s})\bigr{)}^{p},$
we begin with the Malliavin derivative of the product $Y_{T}\,Y_{s}^{-1}\,\sigma(s,X_{s})$.
$D_{t}W_{s}^{p}$ $=D_{t}\Bigl{(}\bigl{(}Y_{T}\,Y_{s}^{-1}\,\sigma(s,X_{s})\bigr{)}^{p}\Bigr{)}.$
  • Case 1: $t\leq s$. In this scenario, a “kick” in the Brownian motion at time $t$ does affect $X_{s}$ (and hence $Y_{s}$). Thus
    $D_{t}\bigl{(}Y_{T}\,Y_{s}^{-1}\,\sigma(s,X_{s})\bigr{)}$ $=D_{t}Y_{T}\;\cdot\;\bigl{(}Y_{s}^{-1}\,\sigma(s,X_{s})\bigr{)}\;+\;Y_{T}\;\,\underbrace{D_{t}\bigl{(}Y_{s}^{-1}\,\sigma(s,X_{s})\bigr{)}}_{\text{chain rule}}.$
    Using Lemma 5.6 we write $D_{t}Y_{T}$ as
    $D_{t}Y_{T}$ $=Z_{T}\,Y_{t}^{-1}\,\sigma(t,X_{t})-Y_{T}\,Y_{t}^{-1}\,Z_{t}\,Y_{t}^{-1}\,\sigma(t,X_{t})$
    $\quad+Y_{T}\,Y_{t}^{-1}\,\partial_{x}\sigma(t,X_{t})\,Y_{t}$
    Note that
    $D_{t}\bigl{(}Y_{s}^{-1}\,\sigma(s,X_{s})\bigr{)}$ $=\bigl{(}D_{t}Y_{s}^{-1}\bigr{)}\,\sigma(s,X_{s})\;+\;Y_{s}^{-1}\,\bigl{(}D_{t}\sigma(s,X_{s})\bigr{)}.$ (5.4)
    For $t\leq s$, the Malliavin derivative of the inverse is
    $D_{t}Y_{s}^{-1}=-Y_{s}^{-1}\big{[}$ $Z_{s}Y_{t}^{-1}\sigma(t,X_{t})$
    $-Y_{s}Y_{t}^{-1}Z_{t}Y_{t}^{-1}\sigma(t,X_{t})$
    $+Y_{s}Y_{t}^{-1}\partial_{x}\sigma(t,X_{t})Y_{t}\big{]}Y_{s}^{-1}.$
    This expression accounts for the second variation processes $Z_{s}$ and $Z_{t}$, as well as the derivative of the diffusion coefficient. Applying the chain rule
    $D_{t}\sigma(s,X_{s})\;=\;\partial_{x}\sigma(s,X_{s})\,\bigl{(}Y_{s}\,Y_{t}^{-1}\,\sigma(t,X_{t})\bigr{)}.$
    Substituting into (5.4) yields
    $D_{t}\bigl{(}Y_{s}^{-1}\,\sigma(s,X_{s})\bigr{)}$ $=-Y_{s}^{-1}\Bigl{[}Z_{s}Y_{t}^{-1}\sigma(t,X_{t})$
    $\quad-Y_{s}Y_{t}^{-1}Z_{t}Y_{t}^{-1}\sigma(t,X_{t})$
    $\quad+Y_{s}Y_{t}^{-1}\partial_{x}\sigma(t,X_{t})Y_{t}\Bigr{]}Y_{s}^{-1}\sigma(s,X_{s})$
    $\quad+Y_{s}^{-1}\partial_{x}\sigma(s,X_{s})\bigl{(}Y_{s}Y_{t}^{-1}\sigma(t,X_{t})\bigr{)}.$
Thus, for $t\leq s$
$\begin{aligned} D_{t}W_{s}^{p}=\Biggl{[}&\Bigl{(}Z_{T}\,Y_{t}^{-1}\,\sigma(t,X_{t})-\,Y_{T}\,Y_{t}^{-1}\,Z_{t}\,Y_{t}^{-1}\,\sigma(t,X_{t})\\ &\quad+\,Y_{T}\,Y_{t}^{-1}\,\partial_{x}\sigma(t,X_{t})\,Y_{t}\Bigr{)}\,Y_{s}^{-1}\,\sigma(s,X_{s})\\ &\quad+\,Y_{T}\Biggl{(}-\,Y_{s}^{-1}\,\Bigl{[}Z_{s}Y_{t}^{-1}\,\sigma(t,X_{t})-\,Y_{s}Y_{t}^{-1}\,Z_{t}Y_{t}^{-1}\,\sigma(t,X_{t})\\ &\quad+\,Y_{s}Y_{t}^{-1}\,\partial_{x}\sigma(t,X_{t})Y_{t}\Bigr{]}\,Y_{s}^{-1}\,\sigma(s,X_{s})\\ &\quad+\,Y_{s}^{-1}\,\partial_{x}\sigma(s,X_{s})\Bigl{(}Y_{s}Y_{t}^{-1}\,\sigma(t,X_{t})\Bigr{)}\Biggr{)}\Biggr{]}^{p}.\end{aligned}$
We place the entire sum inside brackets $\dots]^{p}$ because we are taking the $p$-th component of the resulting vector.
  • Case 2: $t>s$. In this case, a Brownian perturbation at time $t$ does not affect $X_{s}$ (nor $Y_{s}$) because $s<t$. Hence
    $D_{t}\bigl{(}Y_{s}^{-1}\,\sigma(s,X_{s})\bigr{)}$ $=0,$
    and the only contribution is from $D_{t}Y_{T}$. Therefore,
    $D_{t}W_{s}^{p}=\Biggl{[}\Bigl{(}Z_{T}\,Y_{t}^{-1}\,\sigma(t,X_{t})\\ -\,Y_{T}\,Y_{t}^{-1}\,Z_{t}\,Y_{t}^{-1}\,\sigma(t,X_{t})+\,Y_{T}\,Y_{t}^{-1}\,\partial_{x}\sigma(t,X_{t})\,Y_{t}\Bigr{)}\,Y_{s}^{-1}\,\sigma(s,X_{s})\Biggr{]}^{p}.$
Next, we recall that
$\gamma_{X_{T}}^{p,q}=\int_{0}^{T}[Y_{T}Y_{s}^{-1}\sigma(s,X_{s})]^{p}\,\cdot\,[Y_{T}Y_{s}^{-1}\sigma(s,X_{s})]^{q}\,ds.$
Recall that $\gamma_{X_{T}}^{p,q}=\int_{0}^{T}[Y_{T}Y_{s}^{-1}\sigma(s,X_{s})]^{p}\cdot[Y_{T}Y_{s}^{-1}\sigma(s,X_{s})]^{q}\,ds$. Taking the Malliavin derivative and using the product rule
$D_{t}\gamma_{X_{T}}^{p,q}$ $=\int_{0}^{T}\Big{[}D_{t}W_{s}^{p}\cdot W_{s}^{q}+W_{s}^{p}\cdot D_{t}W_{s}^{q}\Big{]}\,ds,$ (5.5)
where $W_{s}^{p}=(Y_{T}Y_{s}^{-1}\sigma(s,X_{s}))^{p}$ and we computed $D_{t}W_{s}^{p}$ in Cases 1 and 2 above. To evaluate (5.5), we split the integration region into $[0,t]$ and $[t,T]$ to reflect the piecewise definitions of $D_{t}W_{s}^{p}$
  • For $s\in[0,t]$: Use the Case 1 expression for $D_{t}W_{s}^{p}$ and $D_{t}W_{s}^{q}$
  • For $s\in[t,T]$: Use the Case 2 expression for $D_{t}W_{s}^{p}$ and $D_{t}W_{s}^{q}$
The resulting expression for $D_{t}\gamma_{X_{T}}^{p,q}$ involves the second variation processes $Z_{t},Z_{s},Z_{T}$ and partial derivatives of the diffusion coefficients, as detailed in the computations above. We now handle the correction term $\sum_{j=1}^{m}\partial_{j}u_{t}(F_{k})\,\cdot\,D_{t}F_{k}^{j}$ that appears in the Skorokhod integral decomposition $\delta(u_{k})=\left.\int_{0}^{T}u_{t}(x)\cdot dB_{t}\right|_{x=F_{k}}\;-\;\sum_{j}\int_{0}^{T}\partial_{j}u_{t}(F_{k})\,\cdot\,D_{t}F_{k}^{j}\,dt$. Hereafter we show how each step follows from the chain rule in Malliavin calculus, the use of $D_{t}(\gamma_{X_{T}}^{-1})=-\,\gamma_{X_{T}}^{-1}\,(D_{t}\gamma_{X_{T}})\,\gamma_{X_{T}}^{-1},$ and the expression for $D_{t}\gamma_{X_{T}}^{p,q}$. We also show how to integrate the resulting expression over $t$. Let us begin by recalling the general formula for the Skorokhod integral
$\delta(u_{k})$ $=\left.\int_{0}^{T}u_{t}(x)\cdot dB_{t}\right|_{x=F_{k}}-\int_{0}^{T}\sum_{j=1}^{m}\partial_{j}u_{t}(F_{k})\,\cdot\,D_{t}\bigl{(}F_{k}^{j}\bigr{)}\,dt.$
The term $\sum_{j=1}^{m}\partial_{j}u_{t}(F_{k})\,\cdot\,D_{t}F_{k}^{j}$ is often called the “correction term.” We have already found that $\partial_{j}u_{t}(F_{k})=\bigl{[}Y_{t}^{-1}\,\sigma(t,X_{t})\bigr{]}_{j},$ and $F_{k}^{j}=e_{j}^{\top}Y_{T}^{\top}\gamma_{X_{T}}^{-1}e_{k}.$
Since $D_{t}(\gamma_{X_{T}}^{-1})=-\,\gamma_{X_{T}}^{-1}\,(D_{t}\gamma_{X_{T}})\,\gamma_{X_{T}}^{-1},$ we obtain
$D_{t}F_{k}^{j}=D_{t}(e_{j}^{\top}Y_{T}^{\top}\gamma_{X_{T}}^{-1}e_{k})=e_{j}^{\top}(D_{t}Y_{T}^{\top})\gamma_{X_{T}}^{-1}e_{k}-e_{j}^{\top}Y_{T}^{\top}\gamma_{X_{T}}^{-1}(D_{t}\gamma_{X_{T}})\gamma_{X_{T}}^{-1}e_{k}.$
Hence,
$\sum_{j=1}^{m}\partial_{j}u_{t}(F_{k})\cdot D_{t}F_{k}^{j}$ $=\sum_{j=1}^{m}\bigl{[}Y_{t}^{-1}\sigma(t,X_{t})\bigr{]}_{j}\left(e_{j}^{\top}(D_{t}Y_{T}^{\top})\gamma_{X_{T}}^{-1}e_{k}-e_{j}^{\top}Y_{T}^{\top}\gamma_{X_{T}}^{-1}(D_{t}\gamma_{X_{T}})\gamma_{X_{T}}^{-1}e_{k}\right).$ (5.6)
Now, recall that $D_{t}\gamma_{X_{T}}^{p,q}$ splits into integrals over $[0,t]$ and $[t,T]$, and contains contributions from terms like $Z_{T}\,Y_{t}^{-1}\,\sigma(t,X_{t})$, $Y_{T}\,Y_{t}^{-1}\,Z_{t}\,Y_{t}^{-1}\,\sigma(t,X_{t})$, $Y_{T}\,Y_{t}^{-1}\,\partial_{x}\sigma(t,X_{t})\,Y_{t},$ and so forth (as detailed in Cases 1 and 2). To complete the calculation of the correction term, we substitute (5.5) into (5.6), reorganise by splitting the integration regions, and integrate from $t=0$ to $t=T$. This yields,
$\int_{0}^{T}\sum_{j=1}^{m}\partial_{j}u_{t}(F_{k})\,\cdot\,D_{t}F_{k}^{j}\,dt$ $=\int_{0}^{T}\sum_{j=1}^{m}\bigl{[}Y_{t}^{-1}\sigma(t,X_{t})\bigr{]}_{j}\cdot A_{jk}(t)\,dt$
$\quad-\int_{0}^{T}\sum_{j=1}^{m}\bigl{[}Y_{t}^{-1}\sigma(t,X_{t})\bigr{]}_{j}\cdot B_{jk}(t)\,dt$
$\quad-\int_{0}^{T}\sum_{j=1}^{m}\bigl{[}Y_{t}^{-1}\sigma(t,X_{t})\bigr{]}_{j}\cdot C_{jk}(t)\,dt$
where
$A_{jk}(t)$ $=e_{j}^{\top}\bigg{[}\sigma(t,X_{t})^{\top}(Y_{t}^{-1})^{\top}Z_{T}^{\top}-\sigma(t,X_{t})^{\top}(Y_{t}^{-1})^{\top}Z_{t}^{\top}(Y_{t}^{-1})^{\top}Y_{T}^{\top}$
$\quad\quad\quad+\left(Y_{T}Y_{t}^{-1}\partial_{x}\sigma(t,X_{t})Y_{t}\right)^{\top}\bigg{]}\gamma_{X_{T}}^{-1}e_{k}$
$B_{jk}(t)$ $=e_{j}^{\top}Y_{T}^{\top}\gamma_{X_{T}}^{-1}\cdot\bigg{[}\int_{0}^{t}I_{1}(t,s)\,ds+\int_{0}^{t}I_{2}(t,s)\,ds\bigg{]}\gamma_{X_{T}}^{-1}e_{k}$
$C_{jk}(t)$ $=e_{j}^{\top}Y_{T}^{\top}\gamma_{X_{T}}^{-1}\cdot\bigg{[}\int_{t}^{T}I_{3}(t,s)\,ds+\int_{t}^{T}I_{4}(t,s)\,ds\bigg{]}\gamma_{X_{T}}^{-1}e_{k}$
with the integrands having components
$I^{p,q}_{1}(t,s)$ $=\bigg{[}\Omega(t)Y_{s}^{-1}\sigma(s,X_{s})\bigg{]}^{p}\cdot[Y_{T}Y_{s}^{-1}\sigma(s,X_{s})]^{q}$
$I^{p,q}_{2}(t,s)$ $=[Y_{T}Y_{s}^{-1}\sigma(s,X_{s})]^{p}\cdot\bigg{[}\Omega(t)Y_{s}^{-1}\sigma(s,X_{s})\bigg{]}^{q}$
$I^{p,q}_{3}(t,s)$ $=\bigg{[}\Omega(t)Y_{s}^{-1}\sigma(s,X_{s})+Y_{T}\Theta(t,s)\bigg{]}^{p}\cdot[Y_{T}Y_{s}^{-1}\sigma(s,X_{s})]^{q}$
$I^{p,q}_{4}(t,s)$ $=[Y_{T}Y_{s}^{-1}\sigma(s,X_{s})]^{p}\cdot\bigg{[}\Omega(t)Y_{s}^{-1}\sigma(s,X_{s})+Y_{T}\Theta(t,s)\bigg{]}^{q},$
where
$\Omega(t)$ $=Z_{T}Y_{t}^{-1}\sigma(t,X_{t})-Y_{T}Y_{t}^{-1}Z_{t}Y_{t}^{-1}\sigma(t,X_{t})+Y_{T}Y_{t}^{-1}\partial_{x}\sigma(t,X_{t})Y_{t}$
$\Theta(t,s)$ $=-Y_{s}^{-1}\bigg{[}Z_{s}Y_{t}^{-1}\sigma(t,X_{t})-Y_{s}Y_{t}^{-1}Z_{t}Y_{t}^{-1}\sigma(t,X_{t})$
$\quad+Y_{s}Y_{t}^{-1}\partial_{x}\sigma(t,X_{t})Y_{t}\bigg{]}Y_{s}^{-1}\sigma(s,X_{s})$
$\quad+Y_{s}^{-1}\partial_{x}\sigma(s,X_{s})\bigg{(}Y_{s}Y_{t}^{-1}\sigma(t,X_{t})\bigg{)}.$
These expansions show how the correction term $\sum_{j=1}^{m}\partial_{j}u_{t}(F_{k})\cdot D_{t}F_{k}^{j}$ expands in terms of $D_{t}\gamma_{X_{T}}^{p,q}$. This yields the final formula for the Skorokhod integral is given in (2.4).
6 State-independent diffusion processes
In this section, we derive a simplified expression for the Skorokhod integral $\delta(u_{k})$ for SDEs with state-independent diffusion coefficients. Such SDEs are of interest in diffusion generative modelling because the noise amplitude remains identical across all sample paths, and can be chosen a priori via an appropriate noise scheduler. From a numerical standpoint, stochastic differential equations with state-independent diffusion coefficients exhibit reduced stiffness compared to those with multiplicative noise. This allows for larger time steps during sampling and offers direct control over the evolution of the marginal variance. This significantly simplifies the computation of likelihood weighting and score-matching losses. As a result, most practical architectures for diffusion-based generative modelling are based on state-independent diffusion coefficients – such as the VP (Variance-Preserving), VE (Variance-Exploding), and sub-VP SDEs introduced in [24]. Hence, consider SDEs of the form
$dX_{t}\;=\;b(t,X_{t})\,dt\;+\;\sigma(t)\,dB_{t},$
where the diffusion coefficient $\sigma:[0,T]\to\mathbb{R}^{m\times d}$ depends on time $t$ only. This property will enable the simplifications of $\delta(u_{k})$ stated in the forthcoming corollary. Also, for the linear case with additive (state-independent) noise, the resulting expression reduces to the classical score formulae for linear Gaussian SDEs, [see [14]].
Corollary 6.1 (Skorokhod integral for state-independent diffusion).
Consider the stochastic differential equation
$dX_{t}=b(t,X_{t})\,dt+\sigma(t)\,dB_{t},$
where the diffusion coefficient $\sigma:[0,T]\to\mathbb{R}^{m\times d}$ depends exclusively on time $t$. Under the assumptions of Theorem 2.1, the Skorokhod integral $\delta(u_{k})$ simplifies to
$\delta(u_{k})$ $=\left.\int_{0}^{T}u_{t}(x)\cdot dB_{t}\right|_{x=F_{k}}$
$\quad-\int_{0}^{T}\sum_{j=1}^{m}\bigl{[}Y_{t}^{-1}\sigma(t)\bigr{]}_{j}\cdot A_{jk}(t)\,dt$
$\quad+\int_{0}^{T}\sum_{j=1}^{m}\bigl{[}Y_{t}^{-1}\sigma(t)\bigr{]}_{j}\cdot B_{jk}(t)\,dt$
$\quad+\int_{0}^{T}\sum_{j=1}^{m}\bigl{[}Y_{t}^{-1}\sigma(t)\bigr{]}_{j}\cdot C_{jk}(t)\,dt$
where
$A_{jk}(t)$ $=e_{j}^{\top}\bigg{[}\sigma(t)^{\top}(Y_{t}^{-1})^{\top}Z_{T}^{\top}-\sigma(t)^{\top}(Y_{t}^{-1})^{\top}Z_{t}^{\top}(Y_{t}^{-1})^{\top}Y_{T}^{\top}\bigg{]}\gamma_{X_{T}}^{-1}e_{k}$
$B_{jk}(t)$ $=e_{j}^{\top}Y_{T}^{\top}\gamma_{X_{T}}^{-1}\cdot\bigg{[}\int_{0}^{t}I_{1}(t,s)\,ds\bigg{]}\gamma_{X_{T}}^{-1}e_{k}$
$C_{jk}(t)$ $=e_{j}^{\top}Y_{T}^{\top}\gamma_{X_{T}}^{-1}\cdot\bigg{[}\int_{t}^{T}I_{2}(t,s)\,ds\bigg{]}\gamma_{X_{T}}^{-1}e_{k}$
with integrands
$I_{1}(t,s)$ $=\bigg{[}\Omega(t)Y_{s}^{-1}\sigma(s)\bigg{]}W_{s}^{\top}+W_{s}\bigg{[}\Omega(t)Y_{s}^{-1}\sigma(s)\bigg{]}^{\top}$
$I_{2}(t,s)$ $=\bigg{[}\Phi(t,s)\bigg{]}W_{s}^{\top}+W_{s}\bigg{[}\Phi(t,s)\bigg{]}^{\top}$
and auxiliary processes
$\Omega(t)$ $=Z_{T}Y_{t}^{-1}\sigma(t)-Y_{T}Y_{t}^{-1}Z_{t}Y_{t}^{-1}\sigma(t)$
$\Phi(t,s)$ $=\Omega(t)Y_{s}^{-1}\sigma(s)-Y_{T}Y_{s}^{-1}\bigg{[}Z_{s}Y_{t}^{-1}\sigma(t)-Y_{s}Y_{t}^{-1}Z_{t}Y_{t}^{-1}\sigma(t)\bigg{]}Y_{s}^{-1}\sigma(s)$
$W_{s}$ $=Y_{T}Y_{s}^{-1}\sigma(s)$
where $u_{t}(x)=x^{\top}Y_{t}^{-1}\sigma(t)$ and $F_{k}=Y_{T}^{\top}\gamma_{X_{T}}^{-1}e_{k}$. The state-independence of $\sigma(t)$ eliminates all terms involving $\partial_{x}\sigma(t,X_{t})$, significantly simplifying the expression compared to the general case.
{funding}
Daniele Venturi was supported by the U.S. Department of Energy (DOE) grant DE-SC0024563.